### v1

# Richard Y. Zhang Dept. of Electrical and Computer Engineering

Abstract
The classical low-rank matrix recovery problem is well-known to exhibit benign nonconvexity under the restricted isometry property (RIP): local optimization is guaranteed to converge to the global optimum, where the ground truth is recovered. We investigate whether benign nonconvexity continues to hold when the factor matrices are constrained to be elementwise nonnegative—a common practical requirement. In the simple setting of a rank-1 nonnegative ground truth, we confirm that benign nonconvexity holds in the fully-observed case with RIP constant $\delta=0$. Surprisingly, however, this property fails to extend to the partially-observed case with any arbitrarily small RIP constant $\delta\to 0^{+}$, irrespective of rank overparameterization. This finding exposes a critical theoretical gap: the continuity argument widely used to explain the empirical robustness of low-rank matrix recovery fundamentally breaks down once nonnegative constraints are imposed.
1 Introduction
Low-rank matrix recovery [14]—fully recovering an unknown low-rank ground truth matrix from partial observations—lies at the heart of machine learning, signal processing, and data analytics. Today, the most common approach is to fit a model matrix $UU^{T}$ (or $LR^{T}$) against the partial observations by locally optimizing the factor matrix $U$ (or $L$ and $R$) using simple gradient-based algorithms. A hallmark of the problem is its benign nonconvexity: under the restricted isometry property (RIP) with a sufficiently small constant $\delta$ [1, 4, 13, 21], local optimization is guaranteed to converge to the globally best possible fit, where the ground truth is fully recovered [1, 4, 13, 21].
However, in many important applications, such as image analysis [8] and clustering [16, 22], imposing elementwise nonnegative constraints $U\geq 0$ (or $L,R\geq 0$) can significantly improve the practical usefulness of the recovered factorization. In a seminal work, Lee and Seung [8] demonstrated how nonnegative matrix factorization automatically learns parts-based decompositions of faces, allowing each learned component (e.g., eyes, nose, mouth) to be interpreted as a physical “part” rather than an abstract algebraic factor. This interpretability arises directly from the nonnegativity constraint, which ensures that the learned components add up in a physically coherent way rather than canceling out due to negative entries. Similarly, nonnegativity constraints can improve low-rank clustering by enforcing disjointedness and preventing “destructive interference” across clusters [16].
Existing gradient-based algorithms for low-rank matrix recovery are easily modified to enforce nonnegativity constraints [10, 22]. In practice, these modified algorithms often perform as well for nonnegative recovery as their unconstrained counterparts do for general recovery. This empirical robustness naturally raises the question: Does benign nonconvexity carry over to the nonnegative setting?
Although nonnegative recovery is NP-hard in the worst case, benign nonconvexity could certainly hold in practically relevant settings where recovery is tractable. As a simple starting point, if the ground truth is both elementwise nonnegative and exactly rank-1, then by the Perron–Frobenius theorem, nonnegative recovery is mathematically equivalent to unconstrained recovery. In this simplified setting, it is reasonable to expect that benign nonconvexity should also transfer to nonnegative recovery.
Indeed, we confirm that benign nonconvexity holds in the fully-observed case with RIP constant $\delta=0$, i.e. when all $n^{2}$ elements of the nonnegative rank-1 ground truth are directly measured. Surprisingly, however, benign nonconvexity does not extend to the partially-observed case with $\delta>0$. Whereas benign nonconvexity holds for unconstrained recovery for any $\delta<1/2$, introducing nonnegative constraints $U\geq 0$ (or $L,R\geq 0$) can create spurious local minima that did not previously exist, even as the RIP constant becomes arbitrarily small $\delta\to 0^{+}$. Figure 1 illustrates this phenomenon for $n=2$ and $r=1$. While the landscape may be benign when considered over its entire domain $\mathbb{R}^{n}$, a slight off-axis rotation induced by partial observations can cause a spurious strict local minimum to appear at the boundary once the domain is restricted to the positive orthant $\mathbb{R}_{+}^{n}$.
Figure 1: Rank-1 counterexample. Left: Contour plot of $f(u)={\textstyle\frac{1}{2}}\|\mathcal{A}(uu^{T}-u_{\star}u_{\star}^{T})\|^{2}$ for $u_{\star}=(1,0)$ over the full domain $\mathbb{R}^{2}$, showing a slight off-axis rotation in the benign optimization landscape. Right: Restricting $u$ to the positive orthant $\mathbb{R}_{+}^{2}$ causes the point $u_{0}=(0,0.5)$ to become a strict local minimum. This same principle is used in 2 to extend the counterexample to all search ranks $r\geq r^{\star},$ all true ranks $r^{\star}\geq 1$, and all matrix orders $n\geq r+r^{\star}$.
In fact, the situation is not improved by increasing the true rank $r^{\star}$ of the ground truth, nor by the overparameterizing the search rank $r$ with respect to $r^{\star}$. No matter how large $r\geq r^{\star}$ is chosen with respect to $r^{\star}\geq 1$, one can always apply the same principle illustrated in Figure 1 to construct counterexamples with arbitrarily small RIP constants $\delta\to 0^{+}$ that admit spurious strict local minima. This stands in contrast to unconstrained recovery, where rank overparameterization was recently shown to improve the optimization landscape, eventually eliminating all spurious local minima [17, 18]. Once nonnegativity constraints are introduced, rank overparameterization does not always improve the optimization landscape.
This negative result highlights a fundamental gap in our understanding of benign nonconvexity: the continuity argument widely used to explain the empirical robustness of low-rank matrix recovery fundamentally breaks down when nonnegative constraints are imposed. As we explain below, benign nonconvexity in the hypothetical fully-observed case ($\delta=0$) is widely used to justify the same behavior in the practical partially-observed case (small $\delta>0$), even in problems like matrix completion that do not fully satisfy RIP. However, for nonnegative recovery, there is a discontinuous transition between benign nonconvexity in the $\delta=0$ case and the existence of spurious local minima for any $\delta\to 0^{+}$. This discontinuity suggests a deeper structural reason for the empirical robustness observed in practice.
1.1 Benign nonconvexity via continuity argument
In classical unconstrained low-rank matrix recovery, benign nonconvexity with partial observations is typically established by appealing to the fully-observed case through a continuity argument. Consider the symmetric low-rank matrix recovery cost
$f:\mathbb{R}^{n\times r}\to\mathbb{R},\quad U\mapsto{\textstyle\frac{1}{2}}\|\mathcal{A}(UU^{T}-U_{\star}U_{\star}^{T})\|^{2},$ (1)
where $U_{\star}\in\mathbb{R}^{n\times r^{\star}}$ is unknown, and the linear operator $\mathcal{A}:\mathbb{R}^{n\times n}\to\mathbb{R}^{m}$ is assumed to satisfy the restricted isometry property with perturbation constant $\delta$:
$(1-\delta)\|M\|^{2}\leq\|\mathcal{A}(M)\|^{2}\leq(1+\delta)\|M\|^{2}\quad\text{ for all }\operatorname{rank}(M)\leq r+r^{\star}.$ (RIP)
[…]
This same mechanism for benign nonconvexity further generalizes to the asymmetric variant. The key idea is to view the $n_{1}\times n_{2}$ matrix $LR^{T}$ as the top-right off-diagonal block of the $n\times n$ matrix $UU^{T}$ of the same rank, where $n=n_{1}+n_{2}$ and $U=[L;R]$ is the column-wise concatenation of the factors. Effectively, we seek to repeat the same proof as the above for the following asymmetric cost function
$g:\mathbb{R}^{n\times r}\to\mathbb{R},\quad[L;R]\mapsto{\textstyle\frac{1}{2}}\|\mathcal{A}(LR^{T}-L_{\star}R_{\star}^{T})\|^{2}.$ (2)
In order to do so, the function $g$ must be augmented with a balancing regularizer
$h:\mathbb{R}^{n\times r}\to\mathbb{R},\quad[L;R]\mapsto{\textstyle\frac{1}{2}}\|L^{T}L-R^{T}R\|^{2},$ (3)
to eliminate the ambiguity in the factorization $LR^{T}=(LG)(RG^{-1})^{T}$. Repeating the same continuity argument yields similar benign nonconvexity guarantees for the asymmetric variant, with only small constant factor adjustments. For example, Zhang’s overparameterization guarantee [18] extends to the asymmetric variant for all $2\delta<(1+\sqrt{r^{\star}/r})^{-1}$.
The same mechanism even extends to “harder” choices of $\mathcal{A}$ that only partially satisfy (RIP). In the matrix completion problem, for example, the condition $\|\mathcal{A}(M)\|^{2}\approx\|M\|^{2}$ holds only when $M$ is both low-rank and sufficiently dense. Nevertheless, Ge et al. [5] established benign nonconvexity using the same continuity argument, by augmenting $f$ with a regularizing term to promote density in the factor matrix $U$. All of these results have been further expanded to when the ground truth matrices $U_{\star},L_{\star},R_{\star}$ are only approximately low-rank, and when only approximate critical points can be computed [4]. We refer the reader to [4, 2] for a complete overview of this proof technique.
1.2 Breakdown of continuity under nonnegativity
It is natural to suspect that the same continuity argument would also carry through to establish benign nonconvexity under nonnegativity. We consider both the symmetric variant
$\min_{U\in\mathbb{R}^{n\times r}}\quad f(U)\quad\text{ subject to }\quad U\geq 0,$ (4)
and the regularized asymmetric variant (for some $\lambda\geq 0$)
$\min_{L\in\mathbb{R}^{n_{1}\times r},R\in\mathbb{R}^{n_{2}\times r}}\quad g([L;R])+\lambda\;h([L;R])\quad\text{ subject to }\quad L\geq 0,\quad R\geq 0,$ (5)
where $f,g,h$ are defined as in the unconstrained case in (1), (2), and (3) but with respect to unknown $U_{\star}\in\mathbb{R}_{+}^{n\times r^{\star}},$ $L_{\star}\in\mathbb{R}_{+}^{n_{1}\times r^{\star}},$ and $R_{\star}\in\mathbb{R}_{+}^{n_{2}\times r^{\star}}$.
In this nonnegative setting, the fully-observed $\delta=0$ instances are already NP-hard in general, as they correspond to completely positive factorization [3] and nonnegative matrix factorization [15] respectively. However, specific instances are indeed solvable in polynomial time or even linear time. For example, the rank $r^{\star}=1$ case corresponds to the computation of the Perron–Frobenius eigenvector, which is solved by the famous PageRank algorithm [6], while the rank $r^{\star}=2$ case corresponds to rotating points on the 2D plane [7]. These same instances are also tractable for small $\delta>0$, as one can first solve unconstrained recovery for the ground truths $U_{\star}U_{\star}^{T}$ or $L_{\star}R_{\star}^{T}$, and then separately recover the nonnegative factors $U_{\star},L_{\star},R_{\star}\geq 0$.
Indeed, we confirm in the simple $\delta=0$ and $r^{\star}=1$ case that both the symmetric variant (4) and the asymmetric variant (5) have no spurious local minima.
Proposition 1.
Let $\mathcal{A}$ satisfy (RIP) with constant $\delta=0$. For $r\geq r^{\star}=1$, the problems (4) and (5) with $\lambda=1/4$ have no spurious local minima.
But in contrast to the unconstrained regime, the same conclusions do not carry over to any value of $\delta>0$, irrespective of how much the search rank $r$ is overparameterized beyond the true rank $r^{\star}$. Below, $\mathbf{1}_{r}$ is the length-$r$ vector of ones, and $I_{r}$ is the $r\times r$ identity matrix.
Theorem 2.
For $r\geq r^{\star}\geq 1$, choose $Q_{1}\in\mathbb{R}_{+}^{n\times r^{\star}}$ and $Q_{2}\in\mathbb{R}_{+}^{n\times r}$ to be the first $r+r^{\star}\leq n$ columns of an $n\times n$ permutation matrix. For any $\alpha,\epsilon>0$, define $\mathcal{A}:\mathbb{R}^{n\times n}\to\mathbb{R}^{n\times n}$ to implement the following map
$X\mapsto X+\frac{\operatorname{tr}Q_{1}^{T}XQ_{1}}{r^{\star}}\begin{bmatrix}Q_{1}&Q_{2}\end{bmatrix}\begin{bmatrix}(-1+\sqrt{1-\delta^{2}})I_{r^{\star}}&-\epsilon\alpha^{2}\mathbf{1}_{r^{\star}}\mathbf{1}_{r}^{T}\\ -\epsilon\alpha^{2}\mathbf{1}_{r}\mathbf{1}_{r^{\star}}^{T}&\alpha^{2}I_{r}\end{bmatrix}\begin{bmatrix}Q_{1}&Q_{2}\end{bmatrix}^{T}$
with $\delta=\alpha^{2}\sqrt{r/r^{\star}+2\epsilon^{2}r}$. Then, $\mathcal{A}$ satisfies (RIP) with constant $\delta$, but the instance of (4) with $U_{\star}=Q_{1}$ has a spurious strict local minimum at $U_{0}=\alpha Q_{2}$. Moveover, the instance of (5) with $L_{\star}=R_{\star}=Q_{1}$ has a spurious second-order critical point at $L_{0}=R_{0}=\alpha Q_{2}$ for any $\lambda\geq 0$, that is also a strict local minimum if $\lambda>0$.
Therefore, we conclude that the standard continuity argument does not generalize to nonnegative low-rank matrix recovery. Our proof of 2 works by verifying the second-order sufficient conditions for local optimality at the stated spurious points. In the inequality-constrained setting, this conceptionally simple task is in fact NP-hard in general, as it amounts to copositive testing on a submatrix of the Hessian [11]. Therefore, the main difficulty of the proof is actually our careful construction of $\mathcal{A}$ to allow copositive testing to be performed by hand; this turns out to be particularly delicate in the asymmetric case.
1.3 Organization
The remainder of this paper is organized is follows. In 2, we formally define what it means for a point $U$ to be a spurious local minimum, and to state the second-order necessary and sufficient conditions for local optimality. In 3, we prove 2 by verifying the sufficient conditions at the spurious point. In 4, we prove 1 by verifying that the necessary conditions imply global optimality when $\delta=0$ and $r^{\star}=1$. Finally, in 5, we conclude by speculating on possible future directions that might lead to benign nonconvexity guarantees.
2 Preliminaries
2.1 Sufficient conditions for local optimality
Consider minimizing a twice differentiable cost $f:\mathbb{R}_{+}^{n\times r}\to\mathbb{R}$. The point $U_{0}\in\mathbb{R}_{+}^{n\times r}$ is said to be a local minimum if its value is minimum within a feasible neighborhood
$f(U_{0})\leq f(U)\quad\text{for all }U\in\mathbb{R}_{+}^{n\times r}\cap B(U_{0},\epsilon)$
where $B(U_{0},\epsilon)$ denotes the open Euclidean ball of radius $\epsilon$ centered at $U_{0}$. It is further said to be a strict local minimum if the inequality is strict for all $U\neq U_{0}$. In either case, we say that $U_{0}$ is spurious if it is strictly worse than the global minimum $f(U_{0})>f(U_{\star})$.
Our proof verifies local optimality through the second-order sufficient condition. In our setting with simple bound counstraints, the first-order part of the condition reads
$U\in\mathbb{R}_{+}^{n\times r},\quad\nabla f(U)\in\mathbb{R}_{+}^{n\times r},\quad\left\langle\nabla f(U),U\right\rangle=0,$ (6)
while the second-order part additionally requires
$\left\langle\nabla^{2}f(U)[\dot{U}],\dot{U}\right\rangle>0\text{ for all }\dot{U}\neq 0,\;\dot{U}\in\mathscr{C}(U,\nabla f(U)),$ (7)
$\text{where }\mathscr{C}(U,G)=\left\{\dot{U}\in\mathbb{R}^{n\times r}:\begin{array}[]{cc}\dot{U}_{i,j}=0&\text{ for }G_{i,j}>0,\\ \dot{U}_{i,j}\geq 0&\text{ for }U_{i,j}=0.\end{array}\right\}$ (10)
The set $\mathscr{C}(U,G)$ here is called the critical cone. It is a classical result that if $U_{0}$ satisfies (6) and (7) then it is a strict local minimum; see [12, Theorem 12.6]. However, as pointed out by Murty and Kadabi [11], verifying that (7) holds is NP-hard in general, as it amounts to testing whether a submatrix of the Hessian is copositive.
In reverse, every local minimum $U_{0}$ must satisfy the second-order necessary condition, which shares the same the first-order condition (6) but has the following nonstrict second-order part
$\left\langle\nabla^{2}f(U)[\dot{U}],\dot{U}\right\rangle\geq 0\text{ for all }\dot{U}\in\mathscr{C}(U,\nabla f(U)).$ (11)
A point $U_{0}$ that satisfies (6) and (11) is called a second-order critical point; it can be either a local minimum or a higher-order saddle point, see [12, Theorem 12.5]. Our proof of the inexistence of spurious local minima for $\delta=0$ and $r^{\star}=1$ work by using (6) and (11) to imply $UU^{T}=U_{\star}U_{\star}^{T}$.
2.2 Gradients and Hessians
To verify that $f(U)=\frac{1}{2}\|\mathcal{A}(UU^{T}-U_{\star}U_{\star}^{T})\|^{2}$ satisfies the second-order sufficient condition (6) and (7), we will need to evaluate its gradient and its Hessian quadratic form. These are written respectively as
$\nabla f(U)=(S_{U}+S_{U}^{T})U,\;\left\langle\nabla^{2}f(U)[\dot{U}],\dot{U}\right\rangle=\left\langle S_{U}+S_{U}^{T},\dot{U}\dot{U}^{T}\right\rangle+\|\mathcal{A}(U\dot{U}^{T}+\dot{U}U^{T})\|^{2},$ (12)
where $S_{U}=\mathcal{A}^{T}\mathcal{A}(UU^{T}-U_{\star}U_{\star}^{T})$. The gradient and Hessian quadratic form of the function $g(U)=\frac{1}{2}\|\mathcal{A}(LR^{T}-L_{\star}R_{\star}^{T})\|^{2}$ at $U=[L;R]$ and $\dot{U}=[\dot{L};\dot{R}]$ are respectively
$\nabla g(U)=\begin{bmatrix}S_{L,R}R\\ S_{L,R}^{T}L\end{bmatrix},\;\left\langle\nabla^{2}g(U)[\dot{U}],\dot{U}\right\rangle=2\left\langle S_{L,R},\dot{L}\dot{R}^{T}\right\rangle+\|\mathcal{A}(L\dot{R}^{T}+\dot{L}R^{T})\|^{2},$ (13)
where $S_{L,R}=\mathcal{A}^{T}\mathcal{A}(LR^{T}-L_{\star}R_{\star}^{T})$. We will also need to evaluate the Hessian cross terms, for $\dot{U}_{1}=[\dot{L}_{1};\dot{R}_{1}]$ and $\dot{U}_{2}=[\dot{L}_{2};\dot{R}_{2}]$:
$\left\langle\nabla^{2}g(U)[\dot{U}_{1}],\dot{U}_{2}\right\rangle=\left\langle S_{L,R},\dot{L}_{1}\dot{R}_{2}^{T}+\dot{L}_{2}\dot{R}_{1}^{T}\right\rangle+\left\langle\mathcal{A}(L\dot{R}_{1}^{T}+\dot{L}_{1}R^{T}),\mathcal{A}(L\dot{R}_{2}^{T}+\dot{L}_{2}R^{T})\right\rangle.$ (14)
For the balancing regularizer $h([L;R])=\frac{1}{2}\|\mathcal{A}(LR^{T}-L_{\star}R_{\star}^{T})\|^{2}$ at $U=[L;R]$ and $\dot{U}=[\dot{L};\dot{R}]$, these are respectively
$\nabla h(U)=\begin{bmatrix}LD\\ -RD\end{bmatrix},\;\left\langle\nabla^{2}h(U)[\dot{U}],\dot{U}\right\rangle=2\left\langle D,\dot{L}^{T}\dot{L}-\dot{R}^{T}\dot{R}\right\rangle+4\|L^{T}\dot{L}-R^{T}\dot{R}\|^{2},$ (15)
where $D=L^{T}L-R^{T}R$.
3 Counterexamples for $\delta>0$ and any $r\geq r^{\star}\geq 1$
3.1 Symmetric factorization
Without loss of generality, we can set $Q=[Q_{1},Q_{2}]=[\mathbf{e}_{1},\dots,\mathbf{e}_{r+r^{\star}}]$ where $\mathbf{e}_{i}$ is the $i$-th column of $I_{n}$. Our proof of the symmetric part of 2 is based on the following key assumption satisfied by our specific choice of $\mathcal{A}$.
Assumption 1.
For $\alpha,\epsilon>0$, let $\mathcal{A}:\mathbb{R}^{n\times n}\to\mathbb{R}^{m}$ satisfy the following
$\|\mathcal{A}(X)\|^{2}=\|X\|^{2}+2\left\langle B,Q^{T}XQ\right\rangle\left\langle C,Q^{T}XQ\right\rangle\quad\text{for all }X\in\mathbb{R}^{n\times n}$
$\text{where }B=\begin{bmatrix}0&-\epsilon\mathbf{1}_{r^{\star}}\mathbf{1}_{r}^{T}\\ -\epsilon\mathbf{1}_{r}\mathbf{1}_{r^{\star}}^{T}&I_{r}\end{bmatrix},\qquad C=\frac{\alpha^{2}}{r^{\star}}\begin{bmatrix}I_{r^{\star}}&0\\ 0&0\end{bmatrix}.$
We first verify that the assumption implies (RIP) with the claimed constant $\delta$.
Lemma 3.
Suppose that 1 holds. Then, (RIP) holds with constant $\delta=\|B\|\|C\|=\alpha^{2}\sqrt{r/r^{\star}+2\epsilon^{2}r}$.
Proof.
We rewrite 1 as follows
$\|\mathcal{A}(X)\|^{2}=\operatorname{vec}(X)^{T}[I+bc^{T}+bc^{T}]\operatorname{vec}(X)$
where $b=\operatorname{vec}(QBQ^{T})$ and $c=\operatorname{vec}(QCQ^{T})$. Therefore, $\mathcal{A}^{T}\mathcal{A}$ is a rank-2 perturbation of the identity operator, with eigenvalues
$\mathrm{eig}(\mathcal{A}^{T}\mathcal{A})=\{1+b^{T}c\pm\|b\|\|c\|,1,\dots,1\}.$
In our case, $b^{T}c=0$ and $\|b\|=\sqrt{r+2\epsilon^{2}rr^{\star}}$ and $\|c\|=\alpha^{2}/\sqrt{r^{\star}}$. Therefore, (RIP) holds with $\delta=\|b\|\|c\|$. ∎
Next, we prove that that $U_{0}=\alpha Q_{2}$ is a strict local minimum by verifying the first-order necessary condition (6) and the second-order sufficient condition (7).
Lemma 4.
Suppose that 1 holds. Then, at the point $U_{0}=\alpha Q_{2}$, the function $f(U)=\frac{1}{2}\|\mathcal{A}(UU^{T}-Q_{1}Q_{1}^{T})\|^{2}$ satisfies
$\nabla f(U_{0})\geq 0,\quad\left\langle\nabla f(U_{0}),U_{0}\right\rangle=0,$
$\left\langle\nabla^{2}f(U_{0})[\dot{U}],\dot{U}\right\rangle\geq 2\alpha^{2}\|\dot{U}\|^{2}\text{ for all }\dot{U}\in\mathscr{C}(U_{0},\nabla f(U_{0})).$
Proof.
We verify for $E=U_{0}U_{0}^{T}-U_{\star}U_{\star}^{T}$ that we have
$S_{U}=\mathcal{A}^{T}\mathcal{A}(E)$ $=\begin{bmatrix}-I&0&0\\ 0&\alpha^{2}I&0\\ 0&0&0\end{bmatrix}+\alpha^{2}r\begin{bmatrix}\frac{\alpha^{2}}{r^{\star}}I&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix}-\alpha^{2}\begin{bmatrix}0&-\epsilon\mathbf{1}\mathbf{1}^{T}&0\\ -\epsilon\mathbf{1}\mathbf{1}^{T}&I&0\\ 0&0&0\end{bmatrix}$
$=\alpha^{2}\begin{bmatrix}\beta I&\epsilon\mathbf{1}\mathbf{1}^{T}&0\\ \epsilon\mathbf{1}\mathbf{1}^{T}&0&0\\ 0&0&0\end{bmatrix}\qquad\text{ where }\beta=\alpha^{2}r/r^{\star}-\alpha^{-2}.$
To verify (6), we use (12) to compute the gradient at $U_{0}$ to be
$\nabla f(U_{0})=2\alpha^{2}\begin{bmatrix}\beta I&\epsilon\mathbf{1}\mathbf{1}^{T}&0\\ \epsilon\mathbf{1}\mathbf{1}^{T}&0&0\\ 0&0&0\end{bmatrix}\begin{bmatrix}0\\ \alpha I\\ 0\end{bmatrix}=2\alpha^{3}\begin{bmatrix}\epsilon\mathbf{1}\mathbf{1}^{T}\\ 0\\ 0\end{bmatrix},$
and therefore $\nabla f(U_{0})\geq 0$ and $\left\langle\nabla f(U_{0}),U_{0}\right\rangle=0$. To verify (7), we first compute the associated critical cone to be
$\mathscr{C}(U_{0},\nabla f(U_{0}))=\left\{\begin{bmatrix}0\\ Y\\ Z\end{bmatrix}:\begin{array}[]{c}Y\in\mathbb{R}^{r\times r},\;Y_{i,j}\geq 0\text{ for all }i\neq j\\ Z\in\mathbb{R}_{+}^{(n-r-r^{\star})\times r}\end{array}\right\}.$
Therefore, for any $\dot{U}\in\mathscr{C}(U_{0},\nabla f(U_{0}))$, we have
$\left\langle S_{U},\dot{U}\dot{U}^{T}\right\rangle=\left\langle\alpha^{2}\begin{bmatrix}\beta I&\epsilon\mathbf{1}\mathbf{1}^{T}&0\\ \epsilon\mathbf{1}\mathbf{1}^{T}&0&0\\ 0&0&0\end{bmatrix},\begin{bmatrix}0\\ Y\\ Z\end{bmatrix}\begin{bmatrix}0\\ Y\\ Z\end{bmatrix}^{T}\right\rangle=0,$
$\|\mathcal{A}(U_{0}\dot{U}^{T}+\dot{U}U_{0}^{T})\|^{2}=\left\|\mathcal{A}\left(\begin{bmatrix}0&0&0\\ 0&\alpha(Y+Y^{T})&\alpha Z^{T}\\ 0&\alpha Z&0\end{bmatrix}\right)\right\|^{2}=\alpha^{2}\left\|\begin{bmatrix}Y+Y^{T}&Z^{T}\\ Z&0\end{bmatrix}\right\|^{2}.$
Substituting the above into the Hessian quadratic form in (12) yields
$\left\langle\nabla^{2}f(U_{0})[\dot{U}],\dot{U}\right\rangle$ $=2\left\langle S_{U},\dot{U}\dot{U}^{T}\right\rangle+\|\mathcal{A}(U_{0}\dot{U}^{T}+\dot{U}U_{0}^{T})\|^{2}$
$=0+2\alpha^{2}\left(\|Y\|^{2}+\left\langle Y,Y^{T}\right\rangle+\|Z\|^{2}\right)$
$\geq 2\alpha^{2}(\|Y\|^{2}+\|Z\|^{2})=2\alpha^{2}\|\dot{U}\|^{2}.$
The second line substituted $\|Y+Y^{T}\|^{2}=\|Y\|^{2}+\|Y^{T}\|^{2}+2\left\langle Y,Y^{T}\right\rangle$. The final line recalls that $Y_{i,j}\geq 0$ for all $i\neq j$, and therefore
$\left\langle Y,Y^{T}\right\rangle=\sum_{i,j}Y_{i,j}Y_{j,i}=\sum_{i}Y_{i,i}^{2}+\sum_{i\neq j}Y_{i,j}Y_{j,i}\geq 0.$
Finally, we verify the specific counterexample stated in the theorem does indeed satisfy our key assumption.
Lemma 5.
The $\mathcal{A}$ defined in 2 satisfies 1.
Proof.
First, we rewrite 1 as follows
$\|\mathcal{A}(X)\|^{2}=\operatorname{vec}(X)^{T}[I+bc^{T}+bc^{T}]\operatorname{vec}(X)$
where $b=\operatorname{vec}(QBQ^{T})$ and $c=\operatorname{vec}(QCQ^{T})$. Next, we verify that the $\mathcal{A}$ definied in 2 satisfies
$\operatorname{vec}(\mathcal{A}(X))=[I+(b+\gamma c)c^{T}]\operatorname{vec}(X)\qquad\text{ where }\gamma=\frac{-1+\sqrt{1-\|b\|^{2}\|c\|^{2}}}{\|c\|^{2}}.$
Here, we recall from 3 that $\delta=\|b\|\|c\|$. Finally, we take the difference
$0=$ $(I+bc^{T}+\gamma cc^{T})^{T}(I+bc^{T}+\gamma cc^{T})-(I+bc^{T}+cb^{T})$
$=$ $2\gamma cc^{T}+(bc^{T}+\gamma cc^{T})^{T}(bc^{T}+\gamma cc^{T})$
$=$ $(\|c\|^{2}\gamma^{2}+2\gamma+\|b\|^{2})cc^{T}=0$
The final line uses $b^{T}c=0$, and then substituted our stated $\gamma$ as a solution to the quadratic equation. ∎
3.2 Asymmetric factorization
Our proof of the asymmetric part of 2 is based on that insight that, for a fixed linear operator $\mathcal{A}$ and target matrix $M_{\star}=U_{\star}U_{\star}^{T}$, it is possible for the asymmetric factorization $LR^{T}$ to inherit a local minimum from the symmetric factorization $UU^{T}$. In other words, if $U=U_{0}$ is a local minimum for $UU^{T}$, then $L=R=U_{0}$ is a local minimum for $LR^{T}$. In order for this inheritance to be possible, the operator $\mathcal{A}$ must satisfy the following key property.
Assumption 2.
Let $\mathcal{A}:\mathbb{R}^{n\times n}\to\mathbb{R}^{m}$ satisfy $\mathcal{A}^{T}\mathcal{A}(E)\in\mathbb{S}^{n}$ for all $E\in\mathbb{S}^{n}$.
In fact, this key property is already implied by our critical assumption from the previous section. Nevertheless, we prefer to isolate 2, because the technique of passing local minima down from $UU^{T}$ to $LR^{T}$ is broadly useful in the study of benign nonconvexity in low-rank factorizations.
Lemma 6.
1 implies 2.
Proof.
1 specifies $\|\mathcal{A}(S+K)\|^{2}=\|\mathcal{A}(S)\|^{2}+\|\mathcal{A}(K)\|^{2}$ for all symmetric $S=S^{T}$ and skew symmetric $K=-K^{T}$. This implies that $\left\langle\mathcal{A}(S),\mathcal{A}(K)\right\rangle=0$ for all $S,K$, which holds only if $\mathcal{A}^{T}\mathcal{A}(S)=[\mathcal{A}^{T}\mathcal{A}(S)]^{T}$ and $\mathcal{A}^{T}\mathcal{A}(K)=-[\mathcal{A}^{T}\mathcal{A}(K)]^{T}$. ∎
Under 2, there is a simple connection between the derivatives of the symmetric $UU^{T}$ and the asymmetric $LR^{T}$ cases. Note that we do not yet impose nonnegativity on the factor matrices $U,L,R$.
Lemma 7.
Suppose that 2 holds. Define the two functions
$f(U)=\|\mathcal{A}(UU^{T}-M^{\star})\|^{2},\quad g([L;R])=\|\mathcal{A}(LR^{T}-M^{\star})\|^{2},$
with the same $M^{\star}$. Then, their gradients are related at $U^{(2)}=[U;U]$ as
$\nabla g(U^{(2)})=\frac{1}{2}[\nabla f(U);\nabla f(U)],$
and their Hessian quadratic forms are related at $U^{(2)}=[U;U]$ for all $\dot{L},\dot{R}$ as
$\left\langle\nabla^{2}g(U^{(2)})[\dot{L};\dot{R}],[\dot{L};\dot{R}]\right\rangle=\left\langle\nabla^{2}f(U)[\dot{U}],\dot{U}\right\rangle-2\left\langle\mathcal{A}(UU^{T}-M^{\star}),\mathcal{A}(\dot{V}\dot{V}^{T})\right\rangle\\ +\|\mathcal{A}(\dot{V}U^{T}-U\dot{V}^{T})\|^{2}$
where $\dot{U}=\frac{1}{2}(\dot{L}+\dot{R})$ and $\dot{V}=\frac{1}{2}(\dot{L}-\dot{R})$.
Proof.
Under 2, the matrix $S=\mathcal{A}^{T}\mathcal{A}(UU^{T}-M^{\star})$ is symmetric, and therefore
$\nabla g(U^{(2)})=\begin{bmatrix}SU\\ S^{T}U\end{bmatrix}=\frac{1}{2}\begin{bmatrix}(S+S^{T})U\\ (S+S^{T})U\end{bmatrix}=\frac{1}{2}\begin{bmatrix}\nabla f(U)\\ \nabla f(U)\end{bmatrix}.$
Next, using (14), we find that the following Hessian cross term is zero,
$\left\langle\nabla^{2}g(U^{(2)})[\begin{bmatrix}\dot{U}\\ \dot{U}\end{bmatrix}],\begin{bmatrix}\dot{V}\\ -\dot{V}\end{bmatrix}\right\rangle$ $=\langle S,\ddot{K}\rangle+\left\langle\mathcal{A}(\dot{U}U^{T}+U\dot{U}^{T}),\mathcal{A}(\dot{K})\right\rangle$
$=0+\left\langle\mathcal{A}^{T}\mathcal{A}(\dot{U}U^{T}+U\dot{U}^{T}),\dot{K}\right\rangle=0,$
because $\ddot{K}=\dot{V}\dot{U}^{T}-\dot{U}\dot{V}^{T}$ and $\dot{K}=\dot{V}U^{T}-U\dot{V}^{T}$ are skew-symmetric, while both $S$ and $\mathcal{A}^{T}\mathcal{A}(\dot{U}U^{T}+U\dot{U}^{T})$ are symmetric under 2. Therefore, the Hessian quadratic form comprises just the two block-diagonal terms, which we can again evaluate using (14):
$\left\langle\nabla^{2}g(U^{(2)})[\begin{bmatrix}\dot{L}\\ \dot{R}\end{bmatrix}],\begin{bmatrix}\dot{L}\\ \dot{R}\end{bmatrix}\right\rangle=\left\langle\nabla^{2}g(U^{(2)})[\begin{bmatrix}\dot{U}\\ \dot{U}\end{bmatrix}],\begin{bmatrix}\dot{U}\\ \dot{U}\end{bmatrix}\right\rangle+\left\langle\nabla^{2}g(U^{(2)})[\begin{bmatrix}\dot{V}\\ -\dot{V}\end{bmatrix}],\begin{bmatrix}\dot{V}\\ -\dot{V}\end{bmatrix}\right\rangle$
$=\left\langle\nabla^{2}f(U)[\dot{U}],\dot{U}\right\rangle+-2\left\langle S,\dot{V}\dot{V}^{T}\right\rangle+\|\mathcal{A}(\dot{V}U^{T}-U\dot{V}^{T})\|^{2}.$
We also recall the derivatives of the balancing regularizer from (15).
Lemma 8.
Let $h([L;R])=\frac{1}{2}\|L^{T}L-R^{T}R\|^{2}$. Its gradient and Hessian quadratic form are evaluated at at $U^{(2)}=[U;U]$ as following for all $\dot{L},\dot{R}$:
$\nabla h(U^{(2)})=0,\quad\left\langle\nabla^{2}h(U^{(2)})[\dot{L};\dot{R}],[\dot{L};\dot{R}]\right\rangle=4\|U^{T}(\dot{L}-\dot{R})\|^{2}.$
We are now ready to verify that the asymmetric case $LR^{T}$ inherits the same spurious local minimum $L=R=U_{0}=\alpha Q_{2}$ from the symmetric case by verifying the first-order necessary condition (6) and the second-order sufficient condition (7) .
Lemma 9.
Suppose that 1 holds. Then, the function
$g_{\lambda}:\mathbb{R}^{n\times r}\to\mathbb{R},\qquad[L;R]\mapsto\frac{1}{2}\|\mathcal{A}(LR^{T}-Q_{1}Q_{1}^{T})\|^{2}+\frac{\lambda}{2}\|L^{T}L-R^{T}R\|^{2}$
with $\lambda\geq 0$ satisfies at the point $U_{0}^{(2)}=\alpha[Q_{2};Q_{2}]$
$\nabla g_{\lambda}(U_{0}^{(2)})\geq 0,\quad\left\langle\nabla g_{\lambda}(U_{0}^{(2)}),U_{0}^{(2)}\right\rangle=0,$
$\left\langle\nabla^{2}g_{\lambda}(U_{0}^{(2)})[\dot{U}^{(2)}],\dot{U}^{(2)}\right\rangle\geq\mu\|\dot{U}^{(2)}\|^{2}\text{ for all }\dot{U}^{(2)}\in\mathscr{C}(U_{0}^{(2)},\nabla g_{\lambda}(U_{0}^{(2)}))$
where $\mu=\min\{4\lambda,1\}\cdot\alpha^{2}$.
Proof.
Let $f(U)=\frac{1}{2}\|\mathcal{A}(UU^{T}-Q_{1}Q_{1}^{T})\|^{2}$. It immediately follows from 4, 7, and 8 that $g_{\lambda}$ satisfies at $U_{0}^{(2)}=[U_{0};U_{0}]$:
$\nabla g_{\lambda}(U_{0}^{(2)})=\frac{1}{2}\begin{bmatrix}\nabla f(U_{0})\\ \nabla f(U_{0})\end{bmatrix}\geq 0,\quad\left\langle\nabla g_{\lambda}(U_{0}^{(2)}),U_{0}^{(2)}\right\rangle=\frac{1}{2}\left\langle\begin{bmatrix}\nabla f(U_{0})\\ \nabla f(U_{0})\end{bmatrix},\begin{bmatrix}U_{0}\\ U_{0}\end{bmatrix}\right\rangle=0.$
Moreover, this shows that the critical cone for $g_{\lambda}$ is just the Cartesian product of two critical cones for $f$:
$\mathscr{C}(U_{0}^{(2)},\nabla g_{\lambda}(U_{0}^{(2)}))=\{[\dot{L};\dot{R}]:\dot{L},\dot{R}\in\mathscr{C}(U_{0},\nabla f(U_{0}))\}.$
For $\dot{L},\dot{R}\in\mathscr{C}(U_{0},\nabla f(U_{0}))$, note that $\dot{U}=\frac{1}{2}(\dot{L}+\dot{R})\in\mathscr{C}(U_{0},\nabla f(U_{0}))$ because it is a cone. However, the same is not necessarily true of $\dot{V}=\frac{1}{2}(\dot{L}-\dot{R})$. Nevertheless, we recall from the proof of 4 that the sparsity patterns of $S=\mathcal{A}^{T}\mathcal{A}(UU^{T}-M^{\star})$ and $\dot{L},\dot{R}$ are disjoint
$S=\alpha^{2}\begin{bmatrix}\beta I&\epsilon 11^{T}&0\\ \epsilon 11^{T}&0&0\\ 0&0&0\end{bmatrix},\quad\dot{L}=\begin{bmatrix}0\\ Y_{L}\\ Z_{L}\end{bmatrix},\quad\dot{R}=\begin{bmatrix}0\\ Y_{R}\\ Z_{R}\end{bmatrix},$
so that $\left\langle S,\dot{V}\dot{V}^{T}\right\rangle=0$. Moreover, writing $Y=Y_{L}-Y_{R}$ and $Z=Z_{L}-Z_{R}$, we have
$\|\mathcal{A}(\dot{V}U_{0}^{T}-U_{0}\dot{V}^{T})\|^{2}=\left\|\mathcal{A}\left(\begin{bmatrix}0&0&0\\ 0&\alpha(Y-Y^{T})&-\alpha Z^{T}\\ 0&\alpha Z&0\end{bmatrix}\right)\right\|^{2}\geq 2\alpha^{2}\|Z\|^{2}.$
The balancing regularizer contributes the following term
$\left\langle\nabla^{2}h(U_{0}^{(2)})[\dot{L};\dot{R}],[\dot{L};\dot{R}]\right\rangle=4\|U_{0}^{T}(\dot{L}-\dot{R})\|^{2}=8\alpha^{2}\|Y\|^{2}.$
Finally, piecing everything together,
$\left\langle\nabla^{2}g_{\lambda}(U^{(2)})[\dot{L};\dot{R}],[\dot{L};\dot{R}]\right\rangle=\underbrace{\left\langle\nabla^{2}f(U_{0})[\dot{U}],\dot{U}\right\rangle}_{\geq 2\alpha^{2}\|\dot{U}\|^{2}}$
$-2\underbrace{\left\langle S,\dot{V}\dot{V}^{T}\right\rangle}_{=0}+\underbrace{\|\mathcal{A}(\dot{V}U_{0}^{T}-U_{0}\dot{V}^{T})\|^{2}}_{\geq 2\alpha^{2}\|Z\|^{2}}+\lambda\underbrace{\left\langle\nabla^{2}h(U_{0}^{(2)})[\dot{L};\dot{R}],[\dot{L};\dot{R}]\right\rangle}_{=8\alpha^{2}\|Y\|^{2}}.$
Substituting $\|\dot{V}\|=\|Y\|^{2}+\|Z\|^{2}$ and $\frac{1}{2}\|[\dot{L};\dot{R}]\|^{2}=\|\dot{U}\|^{2}+\|\dot{V}\|^{2}$ completes the proof. ∎
4 No spurious local minima for $\delta=0$ and $r^{\star}=1$
4.1 Symmetric factorization
To prove that $f(U)=\frac{1}{2}\|UU^{T}-u_{\star}u_{\star}^{T}\|^{2}$ for $u_{\star}\in\mathbb{R}_{+}^{n}$ has no spurious local minima, we first show that every first-order critical point $U$ satisfying
$U\geq 0,\quad\nabla f(U)\geq 0,\quad\left\langle\nabla f(U),U\right\rangle=0$
is either globally optimal, or it is zero. This is clearly implied by the lemma below, since $u_{\star}\geq 0$ and therefore $u_{\star}u_{\star}^{T}U\geq 0$.
Lemma 10.
If $\left\langle\nabla f(U),u_{\star}u_{\star}^{T}U\right\rangle\geq 0$ and $\left\langle\nabla f(U),U\right\rangle\leq 0$, then either $U=0$ or $UU^{T}=u_{\star}u_{\star}^{T}$.
Proof.
Define $v=U^{T}u_{\star}\geq 0$, and observe that
${\textstyle\frac{1}{2}}\left\langle\nabla f(U),u_{\star}v^{T}\right\rangle=\left\langle u_{\star}v^{T},(UU^{T}-u_{\star}u_{\star}^{T})U\right\rangle=\left\langle vv^{T},U^{T}U\right\rangle-\left\langle v^{T}v,u_{\star}^{T}u_{\star}\right\rangle\geq 0,$
and therefore
$\|v\|^{2}\|UU^{T}\|\geq\left\langle vv^{T},U^{T}U\right\rangle\geq\left\langle v^{T}v,u_{\star}^{T}u_{\star}\right\rangle=\|v\|^{2}\|u_{\star}u_{\star}^{T}\|.$
Next, we observe that
$\frac{1}{2}\left\langle\nabla f(U),U\right\rangle=\left\langle U,(UU^{T}-u_{\star}u_{\star}^{T})U\right\rangle=\|UU^{T}\|^{2}-\|v\|^{2}\leq 0.$
If $U\neq 0$, then $\|v\|\geq\|UU^{T}\|\neq 0$, and hence $\|UU^{T}\|\geq\|u_{\star}u_{\star}^{T}\|$. Piecing everything together, we have
$\|UU^{T}-u_{\star}u_{\star}^{T}\|^{2}$ $=\|UU^{T}\|^{2}+\|u_{\star}u_{\star}^{T}\|^{2}-2\|v\|^{2}$
$\leq\|u_{\star}u_{\star}^{T}\|^{2}-\|UU^{T}\|^{2}\leq 0.$
Next, we prove that $U=0$ is never a second-order critical point, except in the trivial case when $u_{\star}=0$. Therefore, we conclude that all second-order critical points of $f$ are always globally optimal.
Lemma 11.
If $u_{\star}\neq 0$, then $\left\langle\nabla^{2}f(0)[\dot{U}],\dot{U}\right\rangle<0$ for some $\dot{U}\in\mathscr{C}(0,\nabla f(0))$.
Proof.
Note that $\nabla f(0)=0$, and therefore $\mathscr{C}(0,\nabla f(0))=\mathbb{R}_{+}^{n\times r}$. Then, $\dot{U}=u_{\star}\mathbf{e}_{1}^{T}\geq 0$ yields $\left\langle\nabla^{2}f(0)[\dot{U}],\dot{U}\right\rangle=2\left\langle-u_{\star}u_{\star}^{T},(u_{\star}\mathbf{e}_{1}^{T})(u_{\star}\mathbf{e}_{1}^{T})^{T}\right\rangle=-2\|u_{\star}\|^{2}\|\dot{U}\|^{2}.$
4.2 Asymmetric factorization
Now we move to the asymmetric case. To prove for $U=[L;R]\in\mathbb{R}_{+}^{n\times}$ and $u_{\star}=[a;b]\in\mathbb{R}_{+}^{n}$ with $\|a\|=\|b\|$ that
$g_{\lambda}(U)=\frac{1}{2}\|LR^{T}-ab^{T}\|^{2}+\frac{\lambda}{2}\|L^{T}L-R^{T}R\|^{2}$
has no spurious local minima when $\lambda=1/4$, we will again show that every first-order critical point $U$ is either globally optimal, or it is zero. We prove this by relating the first-order critical points of $g_{\lambda}$ to first-order critical points of the symmetric function $f(U)=\frac{1}{2}\|UU^{T}-u_{\star}u_{\star}^{T}\|^{2}$, and then evoking 10 for the symmetric case.
Lemma 12.
If $U=[L;R]\geq 0$ satisfies $\nabla g_{\lambda}(U)\geq 0$ and $\left\langle\nabla g_{\lambda}(U),U\right\rangle=0$ for $\lambda=1/4$, then it also satisfies $\left\langle\nabla f(U),u_{\star}u_{\star}^{T}U\right\rangle\geq 0$ and $\left\langle\nabla f(U),U\right\rangle\leq 0$.
Proof.
Let $J=\operatorname{diag}(I_{n_{1}},-I_{n_{2}})$ so that $JU=[L;-R]$. Define the error matrix $E=UU^{T}-u_{\star}u_{\star}^{T}$ and partition it into $2\times 2$ blocks so that $E_{12}=E_{21}^{T}=LR^{T}-ab^{T}$. We observe that, for $\lambda=1/4$, the gradients of $g_{\lambda}$ and $f$ are related
$\nabla g_{\lambda}(U)$ $=\left(\begin{bmatrix}0&E_{12}\\ E_{12}^{T}&0\end{bmatrix}+\frac{1}{2}\begin{bmatrix}L\\ -R\end{bmatrix}\begin{bmatrix}L\\ -R\end{bmatrix}^{T}\right)\begin{bmatrix}L\\ R\end{bmatrix}$
$=\left(\begin{bmatrix}0&E_{12}\\ E_{12}^{T}&0\end{bmatrix}+\frac{1}{2}\begin{bmatrix}E_{11}&-E_{12}\\ -E_{12}^{T}&E_{22}\end{bmatrix}+\frac{1}{2}\begin{bmatrix}a\\ -b\end{bmatrix}\begin{bmatrix}b\\ -b\end{bmatrix}^{T}\right)\begin{bmatrix}L\\ R\end{bmatrix}$
$=\left(\frac{1}{2}\begin{bmatrix}E_{11}&E_{12}\\ E_{12}^{T}&E_{22}\end{bmatrix}+\frac{1}{2}\begin{bmatrix}a\\ -b\end{bmatrix}\begin{bmatrix}b\\ -b\end{bmatrix}^{T}\right)\begin{bmatrix}L\\ R\end{bmatrix}$
$=\frac{1}{4}\nabla f(U)+\frac{1}{2}Ju_{\star}u_{\star}^{T}JU.$
Regarding the extra term, note that $u_{\star}^{T}Ju_{\star}=0$ by definition, and therefore
$\left\langle Ju_{\star}u_{\star}^{T}JU,u_{\star}u_{\star}^{T}U\right\rangle=\left\langle u_{\star}^{T}Ju_{\star},u_{\star}^{T}UU^{T}Ju_{\star}\right\rangle=0,$
$\left\langle Ju_{\star}u_{\star}^{T}JU,U\right\rangle=\|U^{T}Ju_{\star}\|\geq 0.$
Hence, if $\left\langle\nabla g_{\lambda}(U),u_{\star}u_{\star}^{T}U\right\rangle\geq 0$ then $\left\langle\nabla f(U),u_{\star}u_{\star}^{T}U\right\rangle\geq 0$, and if $\left\langle\nabla g_{\lambda}(U),U\right\rangle=0$ then $\left\langle\nabla f(U),U\right\rangle\leq 0$. ∎
Again, we prove that $U=0$ is never a second-order critical point for $g_{\lambda}$, except in the trivial case when $u_{\star}=0$. Therefore, we conclude that all second-order critical points of $g_{\lambda}$ for $\lambda=1/4$ are always globally optimal.
Lemma 13.
If $u_{\star}\neq 0$, then $\left\langle\nabla^{2}g_{\lambda}(0)[\dot{U}],\dot{U}\right\rangle<0$ for some $\dot{U}\in\mathscr{C}(0,\nabla g_{\lambda}(0))$.
Proof.
Note that $\nabla g_{\lambda}(0)=0$, and therefore $\mathscr{C}(0,\nabla g_{\lambda}(0))=\mathbb{R}_{+}^{n\times r}$. Then, $\dot{U}=u_{\star}\mathbf{e}_{1}^{T}\geq 0$ yields $\left\langle\nabla^{2}g_{\lambda}(0)[\dot{U}],\dot{U}\right\rangle=\left\langle\begin{bmatrix}0&-ab^{T}\\ -ba^{T}&0\end{bmatrix},(u_{\star}\mathbf{e}_{1}^{T})(u_{\star}\mathbf{e}_{1}^{T})^{T}\right\rangle=-2\|ab^{T}\|^{2}.$ Finally, since $\|a\|=\|b\|$ holds by assumption, we have $\|ab^{T}\|=\|a\|\|b\|=\frac{1}{2}\|u_{\star}\|^{2}=\frac{1}{2}\|\dot{U}\|^{2}$. ∎
5 Conclusions and future work
This paper presented a counterexample to show that the continuity argument used to establish benign nonconvexity for unconstrained low-rank matrix recovery must fundamentally break down once nonnegative constraints are introduced. This breakdown serves as caution for the structural differences between constrained and unconstrained low-rank optimization; one should not expect results developed in one setting to naively port to the other, even if the empirical behaviors in the two settings appear similar.
Below, we outline some promising directions of future research.
Benign nonconvexity for higher ranks
Although our counterexamples provide spurious local minima for any $\delta>0$ and $r\geq r^{\star}\geq 1$, they do not preclude the possibility that $\delta=0$ might remain benign, assuming of course that the particular instance is tractable in the first place. For example, 1 says that the $\delta=0$ and $r^{\star}=1$ case has no spurious local minima. We believe that the same claim is also true for the $\delta=0$ and $r^{\star}=2$ case, although proving this is much more difficult due to the interaction between the two rank-1 components of the ground truth.
Small-but-positive elements and apocalypses
A key component of our construction in 2 is to induce a spurious local minimum $U_{0}$ that contains both exactly zero elements and also small-but-positive elements. This construction takes advantage of the phase transition between zero and near-zero entries, and is closely reminescent of another recent work on so-called apocalypses in the set of rank-bound matrices [9]. This connection suggests explicitly thresholding the elements of the factor matrix $U,L,R$ as a potential safeguard, that might preserve benign nonconvexity by preventing an algorithm from becoming trapped in shallow but combinatorial regions. Developing such safeguarded algorithms or identifying the conditions under which they restore benign nonconvexity could be a fruitful research direction.
Semidefinite program for constructing counterexamples
In the analysis of benign nonconvexity in unconstrained low-rank matrix recovery, a key proof technique is the use of semidefinite programming for constructing counterexamples that admit spurious local minima [20, 19, 17, 18]. In contrast, our counterexample in 2 for the nonnegative setting was manually constructed by hand. A significant difficulty of formulating a semidefinite program to do this is the need to encode the copositive constraint in the second-order optimality conditions, which renders even the task of verifying them as NP-hard. Nevertheless, a more systematic approach to counterexamples could illuminate deeper principles behind why the benign landscape disappears under nonnegativity constraints.
Acknowledgments
I am grateful to Xiaohui Chen, Salar Fattahi, and Sabrina Zielinski for helpful feedback and insightful suggestions. Financial support for this work was provided by NSF CAREER Award ECCS-2047462 and ONR Award N00014-24-1-2671.