### v2

# Richard Y. Zhang Dept. of Electrical and Computer Engineering

Abstract
Low-rank matrix recovery is well-known to exhibit benign nonconvexity under the restricted isometry property (RIP): every second-order critical point is globally optimal, so local methods provably recover the ground truth. Motivated by the strong empirical performance of projected gradient methods for nonnegative low-rank recovery problems, we investigate whether this benign geometry persists when the factor matrices are constrained to be elementwise nonnegative. In the simple setting of a rank-1 nonnegative ground truth, we confirm that benign nonconvexity holds in the fully-observed case with RIP constant $\delta=0$. This benign nonconvexity, however, is unstable. It fails to extend to the partially-observed case with any arbitrarily small RIP constant $\delta>0$, and to higher-rank ground truths $r^{\star}>1$, regardless of how much the search rank $r\geq r^{\star}$ is overparameterized. Together, these results undermine the standard stability-based explanation for the empirical success of nonconvex methods and suggest that fundamentally different tools are needed to analyze nonnegative low-rank recovery.
1 Introduction
Low-rank matrix recovery [17]—recovering an unknown low-rank matrix from partial observations—is a basic problem in modern machine learning and signal processing. A common approach is to represent the matrix in factored form, such as $UU^{T}$ or $LR^{T}$, and to fit the factored matrix to data using simple gradient-based methods. Although the resulting optimization problem is nonconvex and NP-hard in general, these simple methods often work surprisingly well in practice. A standard explanation is the manifestation of benign nonconvexity in tractable regimes. For example, under the Restricted Isometry Property (RIP) with a sufficiently small constant $\delta$, the optimization landscape becomes devoid of spurious local minima, so local optimization always converges to the global minimum [2, 6, 26].
In clustering [20, 27] and image analysis [9] applications, however, the factors are required to be elementwise nonnegative $U\geq 0$ (or $L,R\geq 0$). Imposing nonnegativity is algorithmically straightforward via a projection step $U_{ij}\leftarrow\max\{U_{ij},0\}$. Empirically, the resulting projected gradient methods often behave much like their unconstrained counterparts: once the underlying problem enters a tractable regime, nonconvex optimization appears to succeed reliably. In particular, a recent nonnegative low-rank method for $K$-means clustering [27] recovers planted Gaussian mixture models from essentially any starting point as soon as the clusters are well separated enough for recovery to be polynomial-time solvable by convex relaxation [4].
This empirical parallel—a sharp transition in optimization behavior coinciding with the onset of tractability—naturally raises the question: Does the benign landscape of RIP low-rank matrix recovery persist once nonnegativity constraints are imposed?
This paper performs a systematic investigation and obtains a largely negative answer. While nonnegative low-rank recovery does exhibit benign nonconvexity in a single base case, this fails to persist under mild extensions that are well-known to be benign in unconstrained recovery:
  • Benign landscape holds with a rank $r^{\star}=1$ ground truth in the fully observed $\delta=0$ regime, when all $n_{1}n_{2}$ elements of the $n_{1}\times n_{2}$ rank-1 matrix are directly measured (Proposition 1).
  • Benign landscape does not extend to the partially observed case $\delta>0$, no matter how small the RIP constant (Theorem 3). As shown in Figure 1, spurious local minima can be induced at the boundary for any $\delta\to 0^{+}$, even though the unconstrained landscape becomes benign once $\delta<1/2$ [25, 21].
  • Benign landscape does not extend to higher-rank ground truths $r^{\star}>1$, even when the specific instance of nonnegative recovery is trivial and easily solved by inspection (Proposition 4).
  • Spurious local minima can persist in the overparameterized regime $r>r^{\star}$, regardless of how large the search rank $r$ is set with respect to the true rank $r^{\star}$ (Theorem 3 and Proposition 4). In contrast, unconstrained recovery enjoys benign landscape as soon as the overparameterization ratio exceeds $r/r^{\star}>[\delta/(1-\delta)]^{2},$ where $\delta$ is the RIP constant at rank $r+r^{\star}$ [21, 22].
As we explain next, these negative results are significant because they undermine the stability argument that is widely used to explain the empirical success of nonconvex optimization algorithms.
Figure 1: Rank-1 counterexample with partial observations. Left: Contour plot of $f(u)={\textstyle\frac{1}{2}}\|\mathcal{A}(uu^{T}-u_{\star}u_{\star}^{T})\|^{2}$ taken from Theorem 3 with $n=2$ and $r=1$. The landscape is benign over the entire domain $\mathbb{R}^{2}$, but the partial observations $\delta>0$ induce a slight off-axis rotation. Right: Restricting $u$ to the positive orthant $\mathbb{R}_{+}^{2}$ causes the boundary point $u_{0}=(0,0.5)$ to become a strict local minimizer.
1.1 Benign landscape via stability argument
A large body of work over the past ten years has established benign nonconvexity across a wide range of nonconvex optimization problems. These results span many instances of unconstrained low-rank matrix recovery, including matrix sensing [2], matrix completion [7], robust PCA [6], and phase retrieval [18], as well as equality-constrained versions, such as phase and orthogonal group synchronization [1]. Despite the diversity of settings, all these works share a common proof strategy. One first identifies a base case “prototype” setting in which the optimization landscape can be analyzed exactly and shown to be benign. On then proves that the landscape remains stable under small changes, such as partial observation, imbalance or asymmetry, higher-rank ground truth, overparameterization, or mild noise.
In unconstrained low-rank matrix recovery, the fully observed symmetric rank-1 problem serves as the most basic prototype. Consider the unconstrained symmetric low-rank matrix recovery cost
$f:\mathbb{R}^{n\times r}\to\mathbb{R},\quad U\mapsto{\textstyle\frac{1}{2}}\|\mathcal{A}(UU^{T}-U_{\star}U_{\star}^{T})\|^{2},$ (1)
where $U_{\star}\in\mathbb{R}^{n\times r^{\star}}$ is unknown, the search rank $r\geq r^{\star}$, and the linear operator $\mathcal{A}:\mathbb{R}^{n\times n}\to\mathbb{R}^{m}$ is assumed to satisfy the restricted isometry property [^1] Throughout this paper, $\|\cdot\|$ denotes the matrix Euclidean (Frobenius) norm.
$(1-\delta)\|M\|^{2}\leq\|\mathcal{A}(M)\|^{2}\leq(1+\delta)\|M\|^{2}\quad\text{ for all }\operatorname{rank}(M)\leq r+r^{\star}.$ (RIP)
When $\delta=0$ and $r=r^{\star}=1$, the benign geometry of $f$ is easily established by verifying that the usual second-order necessary condition is also sufficient for global optimality. This property turns out to be structurally stable; one can show that introducing small partial observation $\delta>0$, increasing the true rank $r^{\star}\geq 1$, or overparameterizing the search rank $r\geq r^{\star}$ does not create new second-order critical points [2, 6].
The same mechanism underlies extensions to asymmetric factorizations. The key idea is to view the $n_{1}\times n_{2}$ matrix $LR^{T}$ as the top-right off-diagonal block of the $n\times n$ matrix $UU^{T}$ of the same rank, where $n=n_{1}+n_{2}$ and $U=[L;R]$ is the column-wise concatenation of the factors. One then analyzes the asymmetric objective
$g:\mathbb{R}^{n\times r}\to\mathbb{R},\quad[L;R]\mapsto{\textstyle\frac{1}{2}}\|\mathcal{A}(LR^{T}-L_{\star}R_{\star}^{T})\|^{2},$ (2)
augmented with the balancing regularizer
$h:\mathbb{R}^{n\times r}\to\mathbb{R},\quad[L;R]\mapsto{\textstyle\frac{1}{2}}\|L^{T}L-R^{T}R\|^{2},$ (3)
which removes the scaling ambiguity $LR^{T}=(LG)(RG^{-T})^{T}$. This reduction expresses the asymmetric problem as a controlled perturbation of the symmetric prototype. The stability of benign landscape in the symmetric case then allows the same property to extend to the asymmetric setting, up to constant-factor adjustments [16].
Finally, the same stability argument extends to more challenging measurement models that only approximately satisfy (RIP). In the matrix completion, for example, the condition $\|\mathcal{A}(M)\|^{2}\approx\|M\|^{2}$ holds only for low-rank matrices that are sufficiently dense. Ge et al. [7] established benign nonconvexity in this setting by augmenting $f$ with a regularizer that enforces density, thereby reducing the problem back to the symmetric RIP prototype. Subsequent work shows that benignness remains stable when the ground truth matrices $U_{\star},L_{\star},R_{\star}$ are only approximately low-rank, and when algorithms converge only to approximate critical points [6].
1.2 Breakdown of landscape stability under nonnegativity
Unfortunately, the same stability argument does not carry through to establish benign nonconvexity under nonnegativity. We consider both the symmetric variant
$\min_{U\in\mathbb{R}^{n\times r}}\quad f(U)\quad\text{ subject to }\quad U\geq 0,$ (4)
and the regularized asymmetric variant (for some $\lambda\geq 0$)
$\min_{L\in\mathbb{R}^{n_{1}\times r},R\in\mathbb{R}^{n_{2}\times r}}\quad g([L;R])+\lambda\;h([L;R])\quad\text{ subject to }\quad L\geq 0,\quad R\geq 0,$ (5)
where $f,g,h$ are defined as in the unconstrained case in (1), (2), and (3), but with respect to $U_{\star}\in\mathbb{R}_{+}^{n\times r^{\star}},$ $L_{\star}\in\mathbb{R}_{+}^{n_{1}\times r^{\star}},$ and $R_{\star}\in\mathbb{R}_{+}^{n_{2}\times r^{\star}}$ that are further assumed to be elementwise nonnegative.
In this nonnegative setting, the fully observed $\delta=0$ problem is NP-hard in general, as it encompasses completely positive factorization [5] and nonnegative matrix factorization [19]. At the same time, many instances are computationally trivial. For example, $r^{\star}=1$ reduces to computing a top eigenvector (or singular vector), and $r^{\star}=2$ reduces to a planar rotation [8]. These instances also remain tractable for small $\delta>0$, since one may first recover $UU^{T}$ or $LR^{T}$ by unconstrained methods and then extract nonnegative factors.
We first confirm that the simplest fully observed rank-1 case is benign.
Proposition 1.
Let $\mathcal{A}$ satisfy (RIP) with constant $\delta=0$. For $r\geq r^{\star}=1$, the problems (4) and (5) with $\lambda=1/4$ have no spurious local minima.
Remark 2.
The choice $\lambda=1/4$ is essential; other values of $\lambda$ would break the proof. This weight matches the balancing regularizer used in prior work on unconstrained low-rank recovery [16, 6].
But in contrast to the unconstrained regime, this benign behavior does not extend to any value of $\delta>0$, regardless of overparameterization. Below, $\mathbf{1}_{r}$ denotes the length-$r$ vector of ones, and $I_{r}$ the $r\times r$ identity matrix.
Theorem 3.
For $r\geq r^{\star}\geq 1$, choose $Q_{1}\in\mathbb{R}_{+}^{n\times r^{\star}}$ and $Q_{2}\in\mathbb{R}_{+}^{n\times r}$ so that the concatenation $\begin{bmatrix}Q_{1}&Q_{2}\end{bmatrix}$ consists of the first $r+r^{\star}\leq n$ columns of an $n\times n$ permutation matrix. For $\epsilon>0$ and $0<\alpha<(r/r^{\star}+2\epsilon^{2}r)^{-1/4}$, define $\mathcal{A}:\mathbb{R}^{n\times n}\to\mathbb{R}^{n\times n}$ to implement the following map
$X\mapsto X+\frac{\operatorname{tr}Q_{1}^{T}XQ_{1}}{r^{\star}}\begin{bmatrix}Q_{1}&Q_{2}\end{bmatrix}\begin{bmatrix}(-1+\sqrt{1-\delta^{2}})I_{r^{\star}}&-\epsilon\alpha^{2}\mathbf{1}_{r^{\star}}\mathbf{1}_{r}^{T}\\ -\epsilon\alpha^{2}\mathbf{1}_{r}\mathbf{1}_{r^{\star}}^{T}&\alpha^{2}I_{r}\end{bmatrix}\begin{bmatrix}Q_{1}&Q_{2}\end{bmatrix}^{T}$
with $\delta=\alpha^{2}\sqrt{r/r^{\star}+2\epsilon^{2}r}$. Then, $\mathcal{A}$ satisfies (RIP) with constant $0<\delta<1$, but the instance of (4) with $U_{\star}=Q_{1}$ has a spurious strict local minimum at $U_{0}=\alpha Q_{2}$. Moreover, the instance of (5) with $L_{\star}=R_{\star}=Q_{1}$ has a spurious second-order critical point at $L_{0}=R_{0}=\alpha Q_{2}$ for any $\lambda\geq 0$, which is a strict local minimum whenever $\lambda>0$.
Thus, the landscape is not stable when moving from the fully observed $\delta=0$ case to arbitrarily small partial observation $\delta>0$. Our proof of Theorem 3 works by verifying the second-order sufficient conditions for local optimality at the stated spurious points. In the inequality-constrained setting, this conceptually simple task is in fact NP-hard in general, as it amounts to copositive testing on a submatrix of the Hessian [12]. Therefore, the main difficulty of the proof is actually our careful construction of $\mathcal{A}$ to allow copositive testing to be performed by hand; this turns out to be particularly delicate in the asymmetric case.
A second failure mode appears even before introducing partial observation. In contrast to unconstrained recovery, where the fully observed case is benign for all true ranks $r^{\star}$, nonnegative recovery can already be non-benign at $\delta=0$ once $r^{\star}>1$. Even the seemingly innocuous $r^{\star}=2$ fully observed setting can exhibit spurious local minima.
Theorem 4.
Let $\mathcal{A}$ satisfy (RIP) with constant $\delta=0$. For $m>r\geq r^{\star}>1$, the instances of the problems (4) and (5) with $U_{\star}=L_{\star}=R_{\star}$ each have a spurious local minimum at $U=L=R=U_{0}$, where
$U_{0}=\begin{bmatrix}\frac{1}{\sqrt{r}}\mathbf{1}_{m}\mathbf{1}_{r}^{T}\\ 0\end{bmatrix},\quad U_{\star}=\begin{bmatrix}\mathbf{1}_{m}&0\\ 0&I_{r^{\star}-1}\end{bmatrix}$.
These examples are computationally trivial. They are instances of the exact factorization problem (for $m>r\geq r^{\star}>1$)
$\text{find }U\in\mathbb{R}_{+}^{n\times r}\ \text{such that}\ UU^{T}=M,\;M=\begin{bmatrix}\mathbf{1}_{m}\mathbf{1}_{m}^{T}&0\\ 0&I_{r^{\star}-1}\end{bmatrix}$.
The matrix $M$ is block diagonal with $r^{\star}$ rank-1 blocks. The correct solution is evident by inspection, so the presence of spurious local minima in this example cannot be attributed to computational difficulty. Instead, it arises from the geometry of the nonnegative factorization itself, exacerbated by extreme imbalance among the blocks.
This counterexample also admits a simple clustering interpretation. The matrix $M$ can be viewed as the Gram matrix of $r^{\star}$ groups of points in $\mathbb{R}^{r^{\star}}$, where each group consists of identical points located at a distinct vertex of the standard simplex. Recovering $U$ then amounts to assigning each point to one of these $r^{\star}$ clusters via a one-hot encoding. For example, the $r^{\star}=2$ instance reduces to the problem of partitioning $m$ points at $(1,0)$ and one point at $(0,1)$ into two clusters.
Proposition 4 then carries a cautionary message for clustering applications. Even when clusters are perfectly separated, severe imbalances in cluster size can induce spurious local minima. Empirical successes reported in the clustering literature may therefore rely, implicitly or explicitly, on additional normalization or balance conditions that preclude such degeneracies. For example, the favorable behavior observed in [27] may be aided by the norm constraint $\|U\|^{2}=r^{\star}$ imposed alongside nonnegativity $U\geq 0$, which help prevent the type of collapse exhibited here.
1.3 Organization
The remainder of this paper is organized is follows. In Section 2, we formally define what it means for a point $U$ to be located at a spurious local minimum, and to state the necessary and the second-order sufficient conditions for local optimality. In Section 3, we prove Theorem 3 by verifying the sufficient conditions at the spurious point. In Section 3, we prove Proposition 4 directly. In Section 5, we prove Proposition 1 by verifying that the necessary conditions imply global optimality when $\delta=0$ and $r^{\star}=1$. Finally, in Section 6, we conclude by speculating on possible future directions that might lead to benign nonconvexity guarantees.
2 Preliminaries
2.1 Necessary and sufficient conditions for local optimality
Consider minimizing a twice differentiable cost $f:\mathbb{R}_{+}^{n\times r}\to\mathbb{R}$. The point $U_{0}\in\mathbb{R}_{+}^{n\times r}$ is said to be at a local minimum (resp. is a local minimizer) if its value is minimum within a feasible neighborhood
$f(U_{0})\leq f(U)\quad\text{for all }U\in\mathbb{R}_{+}^{n\times r}\cap B(U_{0},\epsilon)$
where $B(U_{0},\epsilon)$ denotes the open Euclidean ball of radius $\epsilon>0$ centered at $U_{0}$. It is further said to be a strict local minimum if the inequality is strict for all $U\neq U_{0}$. In either case, we say that $U_{0}$ is spurious if it is strictly worse than the global minimum $f(U_{0})>f(U_{\star})$.
Our proof verifies local optimality through the second-order sufficient condition. The first-order part of the condition reads
$U\in\mathbb{R}_{+}^{n\times r},\quad\nabla f(U)\in\mathbb{R}_{+}^{n\times r},\quad\left\langle\nabla f(U),U\right\rangle=0,$ (6)
while the second-order part additionally requires
$\left\langle\nabla^{2}f(U)[\dot{U}],\dot{U}\right\rangle>0\text{ for all }\dot{U}\neq 0,\;\dot{U}\in\mathscr{C}(U,\nabla f(U)),$ (7)
$\text{where }\mathscr{C}(U,G)=\left\{\dot{U}\in\mathbb{R}^{n\times r}:\begin{array}[]{cc}\dot{U}_{i,j}=0&\text{ for }G_{i,j}>0,\\ \dot{U}_{i,j}\geq 0&\text{ for }U_{i,j}=0.\end{array}\right\}$ (10)
The set $\mathscr{C}(U,G)$ here is called the critical cone. It follows from standard Lagrangian theory that if $U_{0}$ satisfies (6) and (7) then it is a strict local minimizer, see [13, Theorem 12.6]. In fact, as clarified by the copositivity-based framework developed in [3, Theorem 1.1], these conditions imply the stronger conclusion that $U_{0}$ is a strong local minimizer, in the sense that a local quadratic growth bound $f(U)\geq f(U_{0})+\frac{\mu}{2}\|U-U_{0}\|^{2}$ holds in a neighborhood of $U_{0}$. However, as pointed out by Murty and Kabadi [12], verifying that (7) holds is NP-hard in general, as it amounts to testing whether a submatrix of the Hessian is strictly copositive.
In reverse, every local minimizer $U_{0}$ must satisfy the second-order necessary condition, which shares the same first-order condition (6) but has the following nonstrict second-order part
$\left\langle\nabla^{2}f(U)[\dot{U}],\dot{U}\right\rangle\geq 0\text{ for all }\dot{U}\in\mathscr{C}(U,\nabla f(U)).$ (11)
A point $U_{0}$ that satisfies (6) and (11) is called a second-order critical point; it can be either a local minimizer or a higher-order saddle point, see [13, Theorem 12.5]. The fact that all local minimizers satisfy (6) and (11) is due to the linearity of the constraint $U\geq 0$, so the tangent and linearization cones coincide, and the Abadie constraint qualification holds. Our proof of the inexistence of spurious local minima for $\delta=0$ and $r^{\star}=1$ work by using (6) and (11) to imply $UU^{T}=U_{\star}U_{\star}^{T}$.
2.2 Gradients and Hessians
To verify that a point $U$ satisfies the second-order sufficient condition (6) and (7), we will need to evaluate the gradient and the Hessian quadratic form of $f(U)=\frac{1}{2}\|\mathcal{A}(UU^{T}-U_{\star}U_{\star}^{T})\|^{2}$, written
$\nabla f(U)=(S_{U}+S_{U}^{T})U,\qquad S_{U}=\mathcal{A}^{*}\mathcal{A}(UU^{T}-U_{\star}U_{\star}^{T}),$ (12)
$\left\langle\nabla^{2}f(U)[\dot{U}],\dot{U}\right\rangle=\left\langle S_{U}+S_{U}^{T},\dot{U}\dot{U}^{T}\right\rangle+\|\mathcal{A}(U\dot{U}^{T}+\dot{U}U^{T})\|^{2}.$
Here, $\mathcal{A}^{*}:\mathbb{R}^{m}\to\mathbb{R}^{n\times n}$ denotes the adjoint map of $\mathcal{A}$. The gradient and Hessian quadratic form of the function $g(U)=\frac{1}{2}\|\mathcal{A}(LR^{T}-L_{\star}R_{\star}^{T})\|^{2}$ at $U=[L;R]$ and $\dot{U}=[\dot{L};\dot{R}]$ are respectively
$\nabla g(U)=\begin{bmatrix}S_{L,R}R\\ S_{L,R}^{T}L\end{bmatrix},\qquad S_{L,R}=\mathcal{A}^{*}\mathcal{A}(LR^{T}-L_{\star}R_{\star}^{T}),$ (13)
$\left\langle\nabla^{2}g(U)[\dot{U}],\dot{U}\right\rangle=2\left\langle S_{L,R},\dot{L}\dot{R}^{T}\right\rangle+\|\mathcal{A}(L\dot{R}^{T}+\dot{L}R^{T})\|^{2}.$
We will also need to evaluate the Hessian cross terms, for $\dot{U}_{1}=[\dot{L}_{1};\dot{R}_{1}]$ and $\dot{U}_{2}=[\dot{L}_{2};\dot{R}_{2}]$:
$\left\langle\nabla^{2}g(U)[\dot{U}_{1}],\dot{U}_{2}\right\rangle=\left\langle S_{L,R},\dot{L}_{1}\dot{R}_{2}^{T}+\dot{L}_{2}\dot{R}_{1}^{T}\right\rangle+\left\langle\mathcal{A}(L\dot{R}_{1}^{T}+\dot{L}_{1}R^{T}),\mathcal{A}(L\dot{R}_{2}^{T}+\dot{L}_{2}R^{T})\right\rangle.$ (14)
For the balancing regularizer $h([L;R])=\frac{1}{2}\|\mathcal{A}(LR^{T}-L_{\star}R_{\star}^{T})\|^{2}$ at $U=[L;R]$ and $\dot{U}=[\dot{L};\dot{R}]$, these are respectively
$\nabla h(U)=\begin{bmatrix}LD\\ -RD\end{bmatrix},\qquad D=L^{T}L-R^{T}R,$ (15)
$\left\langle\nabla^{2}h(U)[\dot{U}],\dot{U}\right\rangle=2\left\langle D,\dot{L}^{T}\dot{L}-\dot{R}^{T}\dot{R}\right\rangle+4\|L^{T}\dot{L}-R^{T}\dot{R}\|^{2}.$
3 Counterexamples for $\delta>0$ and any $r\geq r^{\star}\geq 1$
3.1 Symmetric factorization
Without loss of generality, we can set $Q=[Q_{1},Q_{2}]=[\mathbf{e}_{1},\dots,\mathbf{e}_{r+r^{\star}}]$ where $\mathbf{e}_{i}$ is the $i$-th column of $I_{n}$. Our proof of Theorem 3 relies on the insight that the geometric landscape of $f(U)$ is entirely determined by the kernel operator $\mathcal{H}=\mathcal{A}^{*}\mathcal{A}$, rather than the specific choice of $\mathcal{A}$ itself. In other words, while there are many linear maps $\mathcal{A}$ that satisfy $\mathcal{A}^{*}\mathcal{A}=\mathcal{H}$, the objective value, gradient, and Hessian of $f$ depend only on $\mathcal{H}$; the specific realization $\mathcal{A}$ and its adjoint $\mathcal{A}^{*}$ are never “seen.”
The following is the critical structural property on $\mathcal{H}$ that underpins this paper.
Assumption 1.
For $\alpha,\epsilon>0$, let the kernel operator $\mathcal{H}\equiv\mathcal{A}^{*}\mathcal{A},$ $\mathcal{H}:\mathbb{R}^{n\times n}\to\mathbb{R}^{n\times n}$ satisfy
$\mathcal{H}(X)=X+C\left\langle B,Q^{T}XQ\right\rangle+B\left\langle C,Q^{T}XQ\right\rangle\quad\text{for all }X\in\mathbb{R}^{n\times n}$
$\text{where }B=\begin{bmatrix}0&-\epsilon\mathbf{1}_{r^{\star}}\mathbf{1}_{r}^{T}\\ -\epsilon\mathbf{1}_{r}\mathbf{1}_{r^{\star}}^{T}&I_{r}\end{bmatrix},\qquad C=\frac{\alpha^{2}}{r^{\star}}\begin{bmatrix}I_{r^{\star}}&0\\ 0&0\end{bmatrix}$.
We first verify that that this structural property also implies the restricted isometry property.
Lemma 5.
Suppose that Assumption 1 holds. Then, (RIP) holds with constant $\delta=\|B\|\|C\|=\alpha^{2}\sqrt{r/r^{\star}+2\epsilon^{2}r}$. Additionally, if $\delta<1$, then $\mathcal{H}$ is positive definite.
Proof.
Assumption 1 can be rewritten as $\operatorname{vec}[\mathcal{H}(X)]=[I+bc^{T}+cb^{T}]\operatorname{vec}(X)$ where $b=\operatorname{vec}(QBQ^{T})$ and $c=\operatorname{vec}(QCQ^{T})$. Therefore, $\mathcal{H}$ has eigenvalues
$\mathrm{eig}(\mathcal{H})=\{1+b^{T}c\pm\|b\|\|c\|,1,\dots,1\}.$
In our case, $b^{T}c=0$ and $\|b\|=\sqrt{r+2\epsilon^{2}rr^{\star}}$ and $\|c\|=\alpha^{2}/\sqrt{r^{\star}}$, so (RIP) holds with $\delta=\|b\|\|c\|$. Moveover, if $\delta<1$, then $\mathrm{eig}(\mathcal{H})>0$. ∎
Now, we show that $U_{0}=\alpha Q_{2}$ is a strict local minimizer for any $\mathcal{A}$ whose $\mathcal{H}=\mathcal{A}^{*}\mathcal{A}$ satisfies Assumption 1. Our special choice of $B$ and $C$ in Assumption 1 critically drives the delicate cancellations needed for the second-order sufficient condition (6)-(7) to hold.
Lemma 6.
Suppose that Assumption 1 holds. Then, the function $f(U)=\frac{1}{2}\|\mathcal{A}(UU^{T}-U_{\star}U_{\star})\|^{2}$ with $U_{\star}=Q_{1}$ satisfies at the point $U_{0}=\alpha Q_{2}$:
$f(U_{0})\geq\frac{r^{\star}}{2}(1-\delta^{2}),\quad\nabla f(U_{0})\geq 0,\quad\left\langle\nabla f(U_{0}),U_{0}\right\rangle=0,$
$\left\langle\nabla^{2}f(U_{0})[\dot{U}],\dot{U}\right\rangle\geq\alpha^{2}\|\dot{U}\|^{2}\text{ for all }\dot{U}\in\mathscr{C}(U_{0},\nabla f(U_{0}))$
where $\delta=\alpha^{2}\sqrt{r/r^{\star}+2\epsilon^{2}r}$.
Proof.
We verify for
$E=U_{0}U_{0}^{T}-U_{\star}U_{\star}^{T}=\begin{bmatrix}-I_{r^{\star}}&0&0\\ 0&\alpha^{2}I_{r}&0\\ 0&0&0\end{bmatrix}$,
that we have
$S_{U}=\mathcal{A}^{*}\mathcal{A}(E)$ $=\begin{bmatrix}-I_{r^{\star}}&0&0\\ 0&\alpha^{2}I_{r}&0\\ 0&0&0\end{bmatrix}+\alpha^{2}r\begin{bmatrix}\frac{\alpha^{2}}{r^{\star}}I_{r^{\star}}&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix}-\alpha^{2}\begin{bmatrix}0&-\epsilon\mathbf{1}\mathbf{1}^{T}&0\\ -\epsilon\mathbf{1}\mathbf{1}^{T}&I_{r}&0\\ 0&0&0\end{bmatrix}$
$=\alpha^{2}\begin{bmatrix}-\beta I_{r^{\star}}&\epsilon\mathbf{1}\mathbf{1}^{T}&0\\ \epsilon\mathbf{1}\mathbf{1}^{T}&0&0\\ 0&0&0\end{bmatrix}\qquad\text{ where }\beta=\alpha^{-2}-\alpha^{2}r/r^{\star}.$
Then, the function value at $U_{0}$ is
$f(U_{0})=\frac{1}{2}\|\mathcal{A}(E)\|^{2}=\frac{1}{2}\left\langle S_{U},E\right\rangle=\frac{1}{2}(r^{\star}-\alpha^{4}r)\geq\frac{r^{\star}}{2}(1-\delta^{2}).$
the last inequality substitutes $\delta=\alpha^{2}\sqrt{r/r^{\star}+2\epsilon^{2}r}\geq\alpha^{2}\sqrt{r/r^{\star}}.$ To verify (6), we use (12) to compute the gradient at $U_{0}$ to be
$\nabla f(U_{0})=2\alpha^{2}\begin{bmatrix}\beta I&\epsilon\mathbf{1}\mathbf{1}^{T}&0\\ \epsilon\mathbf{1}\mathbf{1}^{T}&0&0\\ 0&0&0\end{bmatrix}\begin{bmatrix}0\\ \alpha I\\ 0\end{bmatrix}=2\alpha^{3}\begin{bmatrix}\epsilon\mathbf{1}\mathbf{1}^{T}\\ 0\\ 0\end{bmatrix}$,
and therefore $\nabla f(U_{0})\geq 0$ and $\left\langle\nabla f(U_{0}),U_{0}\right\rangle=0$. To verify (7), we first compute the associated critical cone to be
$\mathscr{C}(U_{0},\nabla f(U_{0}))=\left\{\begin{bmatrix}0\\ Y\\ Z\end{bmatrix}:\begin{array}[]{c}Y\in\mathbb{R}^{r\times r},\;Y_{i,j}\geq 0\text{ for all }i\neq j\\ Z\in\mathbb{R}_{+}^{(n-r-r^{\star})\times r}\end{array}\right\}.$
Therefore, for any $\dot{U}\in\mathscr{C}(U_{0},\nabla f(U_{0}))$, we have
$\left\langle S_{U},\dot{U}\dot{U}^{T}\right\rangle=\left\langle\alpha^{2}\begin{bmatrix}\beta I&\epsilon\mathbf{1}\mathbf{1}^{T}&0\\ \epsilon\mathbf{1}\mathbf{1}^{T}&0&0\\ 0&0&0\end{bmatrix},\begin{bmatrix}0\\ Y\\ Z\end{bmatrix}\begin{bmatrix}0\\ Y\\ Z\end{bmatrix}^{T}\right\rangle=0,$
$\|\mathcal{A}(U_{0}\dot{U}^{T}+\dot{U}U_{0}^{T})\|^{2}=\left\|\mathcal{A}\left(\begin{bmatrix}0&0&0\\ 0&\alpha(Y+Y^{T})&\alpha Z^{T}\\ 0&\alpha Z&0\end{bmatrix}\right)\right\|^{2}=\alpha^{2}\left\|\begin{bmatrix}Y+Y^{T}&Z^{T}\\ Z&0\end{bmatrix}\right\|^{2}.$
Substituting the above into the Hessian quadratic form in (12) yields
$\left\langle\nabla^{2}f(U_{0})[\dot{U}],\dot{U}\right\rangle$ $=2\left\langle S_{U},\dot{U}\dot{U}^{T}\right\rangle+\|\mathcal{A}(U_{0}\dot{U}^{T}+\dot{U}U_{0}^{T})\|^{2}$
$=0+2\alpha^{2}\left(\|Y\|^{2}+\left\langle Y,Y^{T}\right\rangle+\|Z\|^{2}\right)$
$\geq 2\alpha^{2}(\|Y\|^{2}+\|Z\|^{2})=2\alpha^{2}\|\dot{U}\|^{2}.$
The second line substituted $\|Y+Y^{T}\|^{2}=\|Y\|^{2}+\|Y^{T}\|^{2}+2\left\langle Y,Y^{T}\right\rangle$. The final line recalls that $Y_{i,j}\geq 0$ for all $i\neq j$, and therefore
$\left\langle Y,Y^{T}\right\rangle=\sum_{i,j}Y_{i,j}Y_{j,i}=\sum_{i}Y_{i,i}^{2}+\sum_{i\neq j}Y_{i,j}Y_{j,i}\geq 0.$
To summarize, suppose that $\mathcal{A}:\mathbb{R}^{n\times n}\to\mathbb{R}^{m}$ satisfies Assumption 1 with $\epsilon>0$ and $0<\alpha<(r/r^{\star}+2\epsilon^{2}r)^{-1/4}$. By Lemma 5, $\mathcal{A}$ also satisfies (RIP) with $\delta=\alpha^{2}\sqrt{r/r^{\star}+2\epsilon^{2}r}<1$. Nevertheless, by Lemma 6, the second-order sufficient condition (6)-(7) holds at a point $U_{0}$ with $f(U_{0})>f(U_{\star})=0$, so $U_{0}$ is indeed a spurious strict local minimizer. Choosing $\alpha\to 0^{+}$ also sets the RIP constant $\delta\to 0^{+}$, so we conclude that there is no $\delta>0$ that would eliminate all spurious local minima.
Formally, a valid choice of $\mathcal{A}$ as describe above is already guaranteed to exist, because $\mathcal{H}$ is positive definite by Lemma 5, so a decomposition $\mathcal{H}=\mathcal{A}^{*}\mathcal{A}$ always exists. Nevertheless, to provide a transparent, computationally reproducible example (as used to generate Figure 1), Theorem 3 provides a specific realization of $\mathcal{A}$. The following lemma confirms this explicit choice is valid.
Lemma 7.
The $\mathcal{A}$ defined in Theorem 3 satisfies Assumption 1.
Proof.
Assumption 1 can be rewritten as $\operatorname{vec}[\mathcal{H}(X)]=[I+bc^{T}+cb^{T}]\operatorname{vec}(X)$ where $b=\operatorname{vec}(QBQ^{T})$ and $c=\operatorname{vec}(QCQ^{T})$. The $\mathcal{A}$ defined in Theorem 3 satisfies in turn
$\operatorname{vec}(\mathcal{A}(X))=[I+(b+\gamma c)c^{T}]\operatorname{vec}(X)\qquad\text{ where }\gamma=\frac{-1+\sqrt{1-\|b\|^{2}\|c\|^{2}}}{\|c\|^{2}}.$
Here, we recall from Lemma 5 that $\delta=\|b\|\|c\|$. Finally, we take the difference
$(I+bc^{T}+\gamma cc^{T})^{T}(I+bc^{T}+\gamma cc^{T})-(I+bc^{T}+cb^{T})$
$=$ $2\gamma cc^{T}+(bc^{T}+\gamma cc^{T})^{T}(bc^{T}+\gamma cc^{T})$
$=$ $(\|c\|^{2}\gamma^{2}+2\gamma+\|b\|^{2})cc^{T}=0$
The final line uses $b^{T}c=0$, and then substituted our stated $\gamma$ as a solution to the quadratic equation. ∎
3.2 Asymmetric factorization
Our proof of the asymmetric part of Theorem 3 is based on the insight that, for a fixed linear operator $\mathcal{A}$ and target matrix $M_{\star}=U_{\star}U_{\star}^{T}$, it is possible for the asymmetric factorization $LR^{T}$ to inherit a local minimum from the symmetric factorization $UU^{T}$. In other words, if $U=U_{0}$ is a local minimizer for $UU^{T}$, then $L=R=U_{0}$ is a local minimizer for $LR^{T}$. In order for this inheritance to be possible, the operator $\mathcal{A}$ must satisfy the following key property.
Assumption 2.
Let $\mathcal{A}:\mathbb{R}^{n\times n}\to\mathbb{R}^{m}$ satisfy $\mathcal{A}^{*}\mathcal{A}(E)\in\mathbb{S}^{n}$ for all $E\in\mathbb{S}^{n}$.
In fact, this key property is already implied by our critical assumption from the previous section. Nevertheless, we prefer to isolate Assumption 2, because the technique of passing local minima down from $UU^{T}$ to $LR^{T}$ is broadly useful in the study of benign nonconvexity in low-rank factorizations.
Lemma 8.
Assumption 1 implies Assumption 2.
Proof.
Assumption 1 specifies $\|\mathcal{A}(S+K)\|^{2}=\|\mathcal{A}(S)\|^{2}+\|\mathcal{A}(K)\|^{2}$ for all symmetric $S=S^{T}$ and skew symmetric $K=-K^{T}$. This implies that $\left\langle\mathcal{A}(S),\mathcal{A}(K)\right\rangle=0$ for all $S,K$, which holds only if $\mathcal{A}^{*}\mathcal{A}(S)=[\mathcal{A}^{*}\mathcal{A}(S)]^{T}$ and $\mathcal{A}^{*}\mathcal{A}(K)=-[\mathcal{A}^{*}\mathcal{A}(K)]^{T}$. ∎
Under Assumption 2, there is a simple connection between the derivatives of the symmetric $UU^{T}$ and the asymmetric $LR^{T}$ cases. Note that we do not yet impose nonnegativity on the factor matrices $U,L,R$.
Lemma 9.
Suppose that Assumption 2 holds. Define the two functions
$f(U)=\frac{1}{2}\|\mathcal{A}(UU^{T}-M^{\star})\|^{2},\quad g([L;R])=\frac{1}{2}\|\mathcal{A}(LR^{T}-M^{\star})\|^{2},$
with the same $M^{\star}$. Then, their gradients are related at $U^{(2)}=[U;U]$ as
$\nabla g(U^{(2)})=\frac{1}{2}[\nabla f(U);\nabla f(U)],$
and their Hessian quadratic forms are related at $U^{(2)}=[U;U]$ for all $\dot{L},\dot{R}$ as
$\left\langle\nabla^{2}g(U^{(2)})[\dot{L};\dot{R}],[\dot{L};\dot{R}]\right\rangle$ $=\left\langle\nabla^{2}f(U)[\dot{U}],\dot{U}\right\rangle-2\left\langle\mathcal{A}(UU^{T}-M^{\star}),\mathcal{A}(\dot{V}\dot{V}^{T})\right\rangle$
$\qquad\qquad+\|\mathcal{A}(\dot{V}U^{T}-U\dot{V}^{T})\|^{2}$
where $\dot{U}=\frac{1}{2}(\dot{L}+\dot{R})$ and $\dot{V}=\frac{1}{2}(\dot{L}-\dot{R})$.
Proof.
Under Assumption 2, the matrix $S=\mathcal{A}^{*}\mathcal{A}(UU^{T}-M^{\star})$ is symmetric, and therefore
$\nabla g(U^{(2)})=\begin{bmatrix}SU\\ S^{T}U\end{bmatrix}=\frac{1}{2}\begin{bmatrix}(S+S^{T})U\\ (S+S^{T})U\end{bmatrix}=\frac{1}{2}\begin{bmatrix}\nabla f(U)\\ \nabla f(U)\end{bmatrix}$.
Next, using (14), we find that the following Hessian cross term is zero,
$\left\langle\nabla^{2}g(U^{(2)})[\begin{bmatrix}\dot{U}\\ \dot{U}\end{bmatrix}],\begin{bmatrix}\dot{V}\\ -\dot{V}\end{bmatrix}\right\rangle$ $=\langle S,\ddot{K}\rangle+\left\langle\mathcal{A}(\dot{U}U^{T}+U\dot{U}^{T}),\mathcal{A}(\dot{K})\right\rangle$
$=0+\left\langle\mathcal{A}^{*}\mathcal{A}(\dot{U}U^{T}+U\dot{U}^{T}),\dot{K}\right\rangle=0,$
because $\ddot{K}=\dot{V}\dot{U}^{T}-\dot{U}\dot{V}^{T}$ and $\dot{K}=\dot{V}U^{T}-U\dot{V}^{T}$ are skew-symmetric, while both $S$ and $\mathcal{A}^{*}\mathcal{A}(\dot{U}U^{T}+U\dot{U}^{T})$ are symmetric under Assumption 2. Therefore, the Hessian quadratic form comprises just the two block-diagonal terms, which we can again evaluate using (14):
$\left\langle\nabla^{2}g(U^{(2)})[\begin{bmatrix}\dot{L}\\ \dot{R}\end{bmatrix}],\begin{bmatrix}\dot{L}\\ \dot{R}\end{bmatrix}\right\rangle=\left\langle\nabla^{2}g(U^{(2)})[\begin{bmatrix}\dot{U}\\ \dot{U}\end{bmatrix}],\begin{bmatrix}\dot{U}\\ \dot{U}\end{bmatrix}\right\rangle+\left\langle\nabla^{2}g(U^{(2)})[\begin{bmatrix}\dot{V}\\ -\dot{V}\end{bmatrix}],\begin{bmatrix}\dot{V}\\ -\dot{V}\end{bmatrix}\right\rangle$
$=\left\langle\nabla^{2}f(U)[\dot{U}],\dot{U}\right\rangle+2\left\langle S,-\dot{V}\dot{V}^{T}\right\rangle+\|\mathcal{A}(\dot{V}U^{T}-U\dot{V}^{T})\|^{2}.$
We also recall the derivatives of the balancing regularizer from (15).
Lemma 10.
Let $h([L;R])=\frac{1}{2}\|L^{T}L-R^{T}R\|^{2}$. Its gradient and Hessian quadratic form are evaluated at $U^{(2)}=[U;U]$ as follows for all $\dot{L},\dot{R}$:
$\nabla h(U^{(2)})=0,\quad\left\langle\nabla^{2}h(U^{(2)})[\dot{L};\dot{R}],[\dot{L};\dot{R}]\right\rangle=4\|U^{T}(\dot{L}-\dot{R})\|^{2}.$
We are now ready to verify that the asymmetric case $LR^{T}$ inherits the same spurious local minimizer $L=R=U_{0}=\alpha Q_{2}$ from the symmetric case by verifying the first-order necessary condition (6) and the second-order sufficient condition (7) .
Lemma 11.
Suppose that Assumption 1 holds. Then, the function
$g_{\lambda}:\mathbb{R}^{n\times r}\to\mathbb{R},\qquad[L;R]\mapsto\frac{1}{2}\|\mathcal{A}(LR^{T}-Q_{1}Q_{1}^{T})\|^{2}+\frac{\lambda}{2}\|L^{T}L-R^{T}R\|^{2}$
with $\lambda\geq 0$ satisfies at the point $U_{0}^{(2)}=\alpha[Q_{2};Q_{2}]$
$g_{\lambda}(U_{0}^{(2)})\geq\frac{r^{\star}}{2}(1-\delta^{2}),\quad\nabla g_{\lambda}(U_{0}^{(2)})\geq 0,\quad\left\langle\nabla g_{\lambda}(U_{0}^{(2)}),U_{0}^{(2)}\right\rangle=0,$
$\left\langle\nabla^{2}g_{\lambda}(U_{0}^{(2)})[\dot{U}^{(2)}],\dot{U}^{(2)}\right\rangle\geq\mu\|\dot{U}^{(2)}\|^{2}\text{ for all }\dot{U}^{(2)}\in\mathscr{C}(U_{0}^{(2)},\nabla g_{\lambda}(U_{0}^{(2)}))$
where $\delta=\alpha^{2}\sqrt{r/r^{\star}+2\epsilon^{2}r}$ and $\mu=\min\{\lambda,\frac{1}{4}\}\cdot\alpha^{2}$.
Proof.
Let $f(U)=\frac{1}{2}\|\mathcal{A}(UU^{T}-Q_{1}Q_{1}^{T})\|^{2}$. It immediately follows from Lemma 6, Lemma 9, and Lemma 10 that $g_{\lambda}$ satisfies at $U_{0}^{(2)}=[U_{0};U_{0}]$:
$g_{\lambda}(U_{0}^{(2)})\geq g_{0}(U_{0}^{(2)})=f_{0}(U_{0})\geq\frac{r^{\star}}{2}(1-\delta^{2}),$
$\nabla g_{\lambda}(U_{0}^{(2)})=\frac{1}{2}\begin{bmatrix}\nabla f(U_{0})\\ \nabla f(U_{0})\end{bmatrix}\geq 0,\quad\left\langle\nabla g_{\lambda}(U_{0}^{(2)}),U_{0}^{(2)}\right\rangle=\frac{1}{2}\left\langle\begin{bmatrix}\nabla f(U_{0})\\ \nabla f(U_{0})\end{bmatrix},\begin{bmatrix}U_{0}\\ U_{0}\end{bmatrix}\right\rangle=0.$
Moreover, this shows that the critical cone for $g_{\lambda}$ is just the Cartesian product of two critical cones for $f$:
$\mathscr{C}(U_{0}^{(2)},\nabla g_{\lambda}(U_{0}^{(2)}))=\{[\dot{L};\dot{R}]:\dot{L},\dot{R}\in\mathscr{C}(U_{0},\nabla f(U_{0}))\}.$
For $\dot{L},\dot{R}\in\mathscr{C}(U_{0},\nabla f(U_{0}))$, note that $\dot{U}=\frac{1}{2}(\dot{L}+\dot{R})\in\mathscr{C}(U_{0},\nabla f(U_{0}))$ because this is a cone. However, the same is not necessarily true of $\dot{V}=\frac{1}{2}(\dot{L}-\dot{R})$. Nevertheless, we recall from the proof of Lemma 6 that the sparsity patterns of $S=\mathcal{A}^{*}\mathcal{A}(UU^{T}-M^{\star})$ and $\dot{L},\dot{R}$ are disjoint:
$S=\alpha^{2}\begin{bmatrix}\beta I&\epsilon 11^{T}&0\\ \epsilon 11^{T}&0&0\\ 0&0&0\end{bmatrix},\quad\dot{L}=\begin{bmatrix}0\\ Y_{L}\\ Z_{L}\end{bmatrix},\quad\dot{R}=\begin{bmatrix}0\\ Y_{R}\\ Z_{R}\end{bmatrix}$,
so that $\left\langle S,\dot{V}\dot{V}^{T}\right\rangle=0$. Moreover, writing $Y=Y_{L}-Y_{R}$ and $Z=Z_{L}-Z_{R}$, we have
$\|\mathcal{A}(\dot{V}U_{0}^{T}-U_{0}\dot{V}^{T})\|^{2}=\left\|\mathcal{A}\left(\begin{bmatrix}0&0&0\\ 0&\alpha(Y-Y^{T})&-\alpha Z^{T}\\ 0&\alpha Z&0\end{bmatrix}\right)\right\|^{2}\geq 2\alpha^{2}\|Z\|^{2}.$
The balancing regularizer contributes the following term
$\left\langle\nabla^{2}h(U_{0}^{(2)})[\dot{L};\dot{R}],[\dot{L};\dot{R}]\right\rangle=4\|U_{0}^{T}(\dot{L}-\dot{R})\|^{2}=8\alpha^{2}\|Y\|^{2}.$
Finally, piecing everything together,
$\left\langle\nabla^{2}g_{\lambda}(U^{(2)})[\dot{L};\dot{R}],[\dot{L};\dot{R}]\right\rangle=\underbrace{\left\langle\nabla^{2}f(U_{0})[\dot{U}],\dot{U}\right\rangle}_{\geq 2\alpha^{2}\|\dot{U}\|^{2}}$
$-2\underbrace{\left\langle S,\dot{V}\dot{V}^{T}\right\rangle}_{=0}+\underbrace{\|\mathcal{A}(\dot{V}U_{0}^{T}-U_{0}\dot{V}^{T})\|^{2}}_{\geq 2\alpha^{2}\|Z\|^{2}}+\lambda\underbrace{\left\langle\nabla^{2}h(U_{0}^{(2)})[\dot{L};\dot{R}],[\dot{L};\dot{R}]\right\rangle}_{=8\alpha^{2}\|Y\|^{2}}.$
Substituting $\frac{1}{4}\|\dot{V}\|^{2}=\|Y\|^{2}+\|Z\|^{2}$ and $\frac{1}{2}\|[\dot{L};\dot{R}]\|^{2}=\|\dot{U}\|^{2}+\|\dot{V}\|^{2}$ completes the proof. ∎
To summarize, given $\epsilon>0$ and $0<\alpha<(r/r^{\star}+2\epsilon^{2}r)^{-1/4}$, Lemma 7 verifies that the $\mathcal{A}:\mathbb{R}^{n\times n}\to\mathbb{R}^{n\times n}$ stated in Theorem 3 satisfies Assumption 1, and Lemma 5 ensures that $\mathcal{A}$ also satisfies (RIP) with $\delta=\alpha^{2}\sqrt{r/r^{\star}+2\epsilon^{2}r}<1$. Nevertheless, by Lemma 11, the second-order necessary condition (6), (11) holds at a point $(L,R)=(U_{0},U_{0})$ with $g_{\lambda}(U_{0},U_{0})>g_{\lambda}(U_{\star},U_{\star})=0$, so this is indeed a spurious second-order critical point. If additionally $\lambda>0$, then the second-order sufficient condition (6)-(7) holds, so $(L,R)=(U_{0},U_{0})$ is also a spurious strict local minimizer. In either case, choosing $\alpha\to 0^{+}$ also sets $\delta\to 0^{+}$, so we conclude that there is no $\delta>0$ that would eliminate all spurious second-order critical points (resp. all spurious local minima).
4 Counterexamples for $\delta=0$ and any $r\geq r^{\star}>1$
We begin with a basic counterexample in the fully observed symmetric setting, which isolates a geometric failure caused solely by the nonnegativity constraint. To ensure that the local minimum identified below is genuinely spurious, we further require $r\geq r^{\star}=k+1$, or else the global minimum of zero may not be attainable at $U=U_{\star}$. This construction also explains why the rank-$1$ case is special; the point $U_{0}$ stated below with $k=0$ is actually globally optimal, and this is consistent with benign behavior for $r^{\star}=1$.
Lemma 12.
Pick $m>r\geq 1$ and $k\geq 1$. The function
$f:\mathbb{R}_{+}^{(m+k)\times r}\to\mathbb{R},\;U\mapsto\frac{1}{2}\|UU^{T}-U_{\star}U_{\star}^{T}\|^{2}$
admits a local minimum at $U_{0}$, where
$U_{0}=\begin{bmatrix}\frac{1}{\sqrt{r}}\mathbf{1}_{m}\mathbf{1}_{r}^{T}\\ 0_{k\times r}\end{bmatrix},\quad U_{\star}=\begin{bmatrix}\mathbf{1}_{m}&0\\ 0&I_{k}\end{bmatrix}$. (16)
Proof.
Partition each $U=\begin{bmatrix}Y\\ Z^{T}\end{bmatrix}$ for $Y\in\mathbb{R}_{+}^{m\times r}$ and $Z\in\mathbb{R}_{+}^{n\times k}$, and observe that
$f(U)-f(U_{0})=$ $\frac{1}{2}\|UU^{T}-U_{\star}U_{\star}^{T}\|^{2}-\frac{1}{2}\|U_{0}U_{0}^{T}-U_{\star}U_{\star}^{T}\|^{2}$
$=$ $\frac{1}{2}\left(\|YY^{T}-\mathbf{1}\mathbf{1}^{T}\|^{2}+2\|YZ\|^{2}+\|Z^{T}Z-I_{k}\|^{2}\right)-\frac{1}{2}\|I_{k}\|^{2}$
$=$ $\frac{1}{2}\left(\|YY^{T}-\mathbf{1}\mathbf{1}^{T}\|^{2}+\|Z^{T}Z-I_{k}\|^{2}\right)+\|YZ\|^{2}-\|Z\|^{2}$
$\geq$ $\|YZ\|^{2}-\|Z\|^{2}.$
At $U=U_{0}$, we have $Y=Y_{0}=\frac{1}{\sqrt{r}}\mathbf{1}_{m}\mathbf{1}_{r}^{T}$, and hence
$\|YZ\|^{2}=\|\frac{1}{\sqrt{r}}\mathbf{1}_{m}\mathbf{1}_{r}^{T}Z\|^{2}=\frac{m}{r}\|\mathbf{1}_{r}^{T}Z\|^{2}$ $=\frac{m}{r}\sum_{i=1}^{k}\|Ze_{i}\|_{\ell_{1}}^{2}\text{ since }Z_{i,j}\geq 0$
$\geq\frac{m}{r}\sum_{i=1}^{k}\|Ze_{i}\|_{\ell_{2}}^{2}=\frac{m}{r}\|Z\|^{2}.$ (17)
Within the neighborhood $\|U-U_{0}\|\leq\rho=\frac{1}{4r\sqrt{m}}$, we have
$\|Y^{T}Y-Y_{0}^{T}Y_{0}\|$ $=\|\mathrm{sym}[(Y-Y_{0})(Y+Y_{0})^{T}]\|\leq\|Y-Y_{0}\|\|Y+Y_{0}\|$
$\leq\rho(2\|Y_{0}\|+\rho)=\rho(2\sqrt{m}+\rho)\leq\frac{1}{r}\quad\text{ for }m,r\geq 1.$ (18)
Combined, the local suboptimality is lower-bounded
$f(U)-f(U_{0})$ $\geq\|YZ\|^{2}-\|Z\|^{2}=\left\langle Y^{T}Y-I_{r},ZZ^{T}\right\rangle$
$\geq\|Y_{0}Z\|^{2}-\|Z\|^{2}-\|Y^{T}Y-Y_{0}^{T}Y_{0}\|\|Z\|^{2}$
$\geq\frac{1}{r}(m-r-1)\|Z\|^{2}\geq 0\quad\text{ for integer }m>r\geq 1,$
so we conclude that $U_{0}$ is a local minimum within this neighborhood. ∎
The argument extends directly to the asymmetric factorization. To keep the notation light, we drop the factor $\frac{1}{2}$ from the definition of $g$, since it plays no role in the local analysis.
Lemma 13.
Pick $m>r\geq 1$ and $k\geq 1$. The function
$g:\mathbb{R}_{+}^{(m+k)\times r}\times\mathbb{R}_{+}^{(m+k)\times r}\to\mathbb{R},\;(L,R)\mapsto\|LR^{T}-U_{\star}U_{\star}^{T}\|^{2}$
admits a local minimum at $L=R=U_{0}$, where $U_{0},U_{\star}$ are the same as (16).
Proof.
Using the same partition of $U$ for both $L$ and $R$, we have $g(U_{0},U_{0})=\|I_{k}\|^{2}$ and
$g(L,R)-g(U_{0},U_{0})=$ $\|LR^{T}-U_{\star}U_{\star}^{T}\|^{2}-\|U_{0}U_{0}^{T}-U_{\star}U_{\star}^{T}\|^{2}$
$=$ $\left(\|Y_{L}Y_{R}^{T}-\mathbf{1}\mathbf{1}^{T}\|^{2}+\|Y_{L}Z_{R}\|^{2}+\|Y_{R}Z_{L}\|^{2}+\|Z_{L}^{T}Z_{R}-I_{k}\|^{2}\right)-\|I_{k}\|^{2}$
$=$ $\left(\|Y_{L}Y_{R}^{T}-\mathbf{1}\mathbf{1}^{T}\|^{2}+\|Z_{L}^{T}Z_{R}\|^{2}\right)+\|Y_{L}Z_{R}\|^{2}+\|Y_{R}Z_{L}\|^{2}-2\left\langle Z_{L},Z_{R}\right\rangle$
$\geq$ $\|Y_{L}Z_{R}\|^{2}+\|Y_{R}Z_{L}\|^{2}-2\left\langle Z_{L},Z_{R}\right\rangle$
$\geq$ $\left(\|Y_{L}Z_{R}\|^{2}-\|Z_{R}\|^{2}\right)+\left(\|Y_{R}Z_{L}\|^{2}-\|Z_{L}\|^{2}\right)$
$=$ $\left\langle Y_{L}^{T}Y_{L}-I,Z_{R}Z_{R}^{T}\right\rangle+\left\langle Y_{R}^{T}Y_{R}-I,Z_{L}Z_{L}^{T}\right\rangle$
The second inequality is because $\|Z_{L}+Z_{R}\|^{2}=\|Z_{L}\|^{2}+\|Z_{R}\|^{2}-2\left\langle Z_{L},Z_{R}\right\rangle\geq 0$. In analogy to (17), at $L=R=U_{0}$, we have $Y_{L}=Y_{R}=Y_{0}=\frac{1}{\sqrt{r}}\mathbf{1}_{m}\mathbf{1}_{r}^{T}$, and hence
$\left\langle Y_{0}^{T}Y_{0}-I,Z_{L}Z_{L}^{T}+Z_{R}Z_{R}^{T}\right\rangle\geq(\frac{m}{r}-1)(\|Z_{L}\|^{2}+\|Z_{R}\|^{2}).$
In analogy to (18), within the neighborhood $\|(L,R)-(U_{0},U_{0})\|\leq\rho=\frac{1}{4r\sqrt{m}}$, we have
$\max\{\|Y_{L}^{T}Y_{L}-Y_{0}^{T}Y_{0}\|,\|Y_{R}^{T}Y_{R}-Y_{0}^{T}Y_{0}\|\}\leq\frac{1}{r}\quad\text{ for }m,r\geq 1.$
Combined, the local suboptimality is lower-bounded
$g(L,R)-g(U_{0},U_{0})$ $\geq\left\langle Y_{L}^{T}Y_{L}-I,Z_{R}Z_{R}^{T}\right\rangle+\left\langle Y_{R}^{T}Y_{R}-I,Z_{L}Z_{L}^{T}\right\rangle$
$\geq(\frac{m}{r}-1-\frac{1}{r})\|Z_{R}\|^{2}+(\frac{m}{r}-1-\frac{1}{r})\|Z_{L}\|^{2}$
$\geq 0\quad\text{ for integer }m>r\geq 1,$
so we conclude that $(L,R)=(U_{0},U_{0})$ is a local minimizer within the neighborhood. ∎
Finally, we show that this spurious local minimum persists after adding the standard balancing regularizer used to eliminate scaling ambiguity in asymmetric factorizations.
Lemma 14.
Let $g_{\lambda}(L,R)=g_{0}(L,R)+\frac{\lambda}{2}\|L^{T}L-R^{T}R\|^{2}$ for $g_{0}:\mathbb{R}_{+}^{n\times r}\times\mathbb{R}_{+}^{n\times r}\to\mathbb{R}$ and $\lambda\geq 0$. If $(L_{0},R_{0})$ with $L_{0}^{T}L_{0}=R_{0}^{T}R_{0}$ is a local minimizer for $g_{0}$, then it is also a local minimizer for $g_{\lambda}$.
Proof.
By definition, there exists a neighborhood $\|(L,R)-(L_{0},R_{0})\|\leq\rho$ such that $g_{0}(L,R)\geq g_{0}(L_{0},R_{0})$. Within that same neighborhood, $g_{\lambda}(L,R)\geq g_{0}(L,R)\geq g_{0}(L_{0},R_{0})=g_{\lambda}(L_{0},R_{0}).$
5 No spurious local minima for $\delta=0$ and $r^{\star}=1$
5.1 Symmetric factorization
To prove that $f(U)=\frac{1}{2}\|UU^{T}-u_{\star}u_{\star}^{T}\|^{2}$ for $u_{\star}\in\mathbb{R}_{+}^{n}$ has no spurious local minima, we first show that every first-order critical point $U$ satisfying
$U\geq 0,\quad\nabla f(U)\geq 0,\quad\left\langle\nabla f(U),U\right\rangle=0$
is either globally optimal, or it is zero. This is clearly implied by the lemma below, since $u_{\star}\geq 0$ and therefore $u_{\star}u_{\star}^{T}U\geq 0$.
Lemma 15.
If $\left\langle\nabla f(U),u_{\star}u_{\star}^{T}U\right\rangle\geq 0$ and $\left\langle\nabla f(U),U\right\rangle\leq 0$, then either $U=0$ or $UU^{T}=u_{\star}u_{\star}^{T}$.
Proof.
Define $v=U^{T}u_{\star}\geq 0$, and observe that
$0\;\leq\;{\textstyle\frac{1}{2}}\left\langle\nabla f(U),u_{\star}v^{T}\right\rangle=\left\langle u_{\star}v^{T},(UU^{T}-u_{\star}u_{\star}^{T})U\right\rangle=\left\langle vv^{T},U^{T}U\right\rangle-\left\langle v^{T}v,u_{\star}^{T}u_{\star}\right\rangle,$
and therefore
$\|v\|^{2}\|UU^{T}\|\geq\left\langle vv^{T},U^{T}U\right\rangle\geq\left\langle v^{T}v,u_{\star}^{T}u_{\star}\right\rangle=\|v\|^{2}\|u_{\star}u_{\star}^{T}\|.$
Next, we observe that
$\frac{1}{2}\left\langle\nabla f(U),U\right\rangle=\left\langle U,(UU^{T}-u_{\star}u_{\star}^{T})U\right\rangle=\|UU^{T}\|^{2}-\|v\|^{2}\leq 0.$
If $U\neq 0$, then $\|v\|\geq\|UU^{T}\|\neq 0$, and hence $\|UU^{T}\|\geq\|u_{\star}u_{\star}^{T}\|$. Piecing everything together, we have
$\|UU^{T}-u_{\star}u_{\star}^{T}\|^{2}$ $=\|UU^{T}\|^{2}+\|u_{\star}u_{\star}^{T}\|^{2}-2\|v\|^{2}$
$\leq\|u_{\star}u_{\star}^{T}\|^{2}-\|UU^{T}\|^{2}\leq 0.$
Next, we prove that $U=0$ is never a second-order critical point, except in the trivial case when $u_{\star}=0$. Therefore, we conclude that all second-order critical points of $f$ are always globally optimal.
Lemma 16.
If $u_{\star}\neq 0$, then $\left\langle\nabla^{2}f(0)[\dot{U}],\dot{U}\right\rangle<0$ for some $\dot{U}\in\mathscr{C}(0,\nabla f(0))$.
Proof.
Note that $\nabla f(0)=0$, and therefore $\mathscr{C}(0,\nabla f(0))=\mathbb{R}_{+}^{n\times r}$. Then, $\dot{U}=u_{\star}\mathbf{e}_{1}^{T}\geq 0$ yields $\left\langle\nabla^{2}f(0)[\dot{U}],\dot{U}\right\rangle=2\left\langle-u_{\star}u_{\star}^{T},(u_{\star}\mathbf{e}_{1}^{T})(u_{\star}\mathbf{e}_{1}^{T})^{T}\right\rangle=-2\|u_{\star}\|^{2}\|\dot{U}\|^{2}.$
5.2 Asymmetric factorization
Now we move to the asymmetric case. To prove for $U=[L;R]\in\mathbb{R}_{+}^{n\times r}$ and $u_{\star}=[a;b]\in\mathbb{R}_{+}^{n}$ with $\|a\|=\|b\|$ that
$g_{\lambda}(U)=\frac{1}{2}\|LR^{T}-ab^{T}\|^{2}+\frac{\lambda}{2}\|L^{T}L-R^{T}R\|^{2}$
has no spurious local minima when $\lambda=1/4$, we will again show that every first-order critical point $U$ is either globally optimal, or it is zero. We prove this by relating the first-order critical points of $g_{\lambda}$ to first-order critical points of the symmetric function $f(U)=\frac{1}{2}\|UU^{T}-u_{\star}u_{\star}^{T}\|^{2}$, and then evoking Lemma 15 for the symmetric case.
Lemma 17.
If $U=[L;R]\geq 0$ satisfies $\nabla g_{\lambda}(U)\geq 0$ and $\left\langle\nabla g_{\lambda}(U),U\right\rangle=0$ for $\lambda=1/4$, then it also satisfies $\left\langle\nabla f(U),u_{\star}u_{\star}^{T}U\right\rangle\geq 0$ and $\left\langle\nabla f(U),U\right\rangle\leq 0$.
Proof.
Let $J=\operatorname{diag}(I_{n_{1}},-I_{n_{2}})$ so that JU=[L;R]JU=[L;-R]. Define the error matrix $E=UU^{T}-u_{\star}u_{\star}^{T}$ and partition it into $2\times 2$ blocks so that $E_{12}=E_{21}^{T}=LR^{T}-ab^{T}$. We observe that, for $\lambda=1/4$, the gradients of $g_{\lambda}$ and $f$ are related
$\nabla g_{\lambda}(U)$ $=\left(\begin{bmatrix}0&E_{12}\\ E_{12}^{T}&0\end{bmatrix}+\frac{1}{2}\begin{bmatrix}L\\ -R\end{bmatrix}\begin{bmatrix}L\\ -R\end{bmatrix}^{T}\right)\begin{bmatrix}L\\ R\end{bmatrix}$
$=\left(\begin{bmatrix}0&E_{12}\\ E_{12}^{T}&0\end{bmatrix}+\frac{1}{2}\begin{bmatrix}E_{11}&-E_{12}\\ -E_{12}^{T}&E_{22}\end{bmatrix}+\frac{1}{2}\begin{bmatrix}a\\ -b\end{bmatrix}\begin{bmatrix}b\\ -b\end{bmatrix}^{T}\right)\begin{bmatrix}L\\ R\end{bmatrix}$
$=\left(\frac{1}{2}\begin{bmatrix}E_{11}&E_{12}\\ E_{12}^{T}&E_{22}\end{bmatrix}+\frac{1}{2}\begin{bmatrix}a\\ -b\end{bmatrix}\begin{bmatrix}b\\ -b\end{bmatrix}^{T}\right)\begin{bmatrix}L\\ R\end{bmatrix}$
$=\frac{1}{4}\nabla f(U)+\frac{1}{2}Ju_{\star}u_{\star}^{T}JU.$
Regarding the extra term, note that $u_{\star}^{T}Ju_{\star}=0$ by definition, and therefore
$\left\langle Ju_{\star}u_{\star}^{T}JU,u_{\star}u_{\star}^{T}U\right\rangle=\left\langle u_{\star}^{T}Ju_{\star},u_{\star}^{T}UU^{T}Ju_{\star}\right\rangle=0,$
$\left\langle Ju_{\star}u_{\star}^{T}JU,U\right\rangle=\|U^{T}Ju_{\star}\|\geq 0.$
Hence, if $\left\langle\nabla g_{\lambda}(U),u_{\star}u_{\star}^{T}U\right\rangle\geq 0$ then $\left\langle\nabla f(U),u_{\star}u_{\star}^{T}U\right\rangle\geq 0$, and if $\left\langle\nabla g_{\lambda}(U),U\right\rangle=0$ then $\left\langle\nabla f(U),U\right\rangle\leq 0$. ∎
Again, we prove that $U=0$ is never a second-order critical point for $g_{\lambda}$, except in the trivial case when $u_{\star}=0$. Therefore, we conclude that all second-order critical points of $g_{\lambda}$ for $\lambda=1/4$ are always globally optimal.
Lemma 18.
If $u_{\star}\neq 0$, then $\left\langle\nabla^{2}g_{\lambda}(0)[\dot{U}],\dot{U}\right\rangle<0$ for some $\dot{U}\in\mathscr{C}(0,\nabla g_{\lambda}(0))$.
Proof.
Note that $\nabla g_{\lambda}(0)=0$, and therefore $\mathscr{C}(0,\nabla g_{\lambda}(0))=\mathbb{R}_{+}^{n\times r}$. Then, $\dot{U}=u_{\star}\mathbf{e}_{1}^{T}\geq 0$ yields $\left\langle\nabla^{2}g_{\lambda}(0)[\dot{U}],\dot{U}\right\rangle=\left\langle\begin{bmatrix}0&-ab^{T}\\ -ba^{T}&0\end{bmatrix},(u_{\star}\mathbf{e}_{1}^{T})(u_{\star}\mathbf{e}_{1}^{T})^{T}\right\rangle=-2\|ab^{T}\|^{2}.$ Finally, since $\|a\|=\|b\|$ holds by assumption, we have $\|ab^{T}\|=\|a\|\|b\|=\frac{1}{2}\|u_{\star}\|^{2}=\frac{1}{2}\|\dot{U}\|^{2}$. ∎
6 Conclusions and future work
This paper shows that the benign nonconvex geometry long associated with RIP-based low-rank recovery is not robust to the introduction of nonnegativity constraints. Although a single fully observed rank-1 case remains benign, this property breaks down under arbitrarily small partial observation, higher-rank ground truths, and overparameterization, with explicit spurious local minima appearing at the boundary. These failures occur in settings that are computationally trivial and easily solvable by inspection, so they must stem from the geometry of nonnegative factorization itself rather than algorithmic hardness. As a result, the standard stability-based explanation that successfully unifies much of unconstrained low-rank recovery does not extend to the nonnegative setting. Understanding the empirical success of projected and constrained methods in nonnegative low-rank recovery applications such as clustering therefore requires new analytical frameworks that go beyond existing stability-based benign landscape arguments.
Below, we outline some promising directions of future research.
Small-but-positive elements and apocalypses
A key component of our construction in Theorem 3 is to induce a spurious local minimizer $U_{0}$ that contains both exactly zero elements and also small-but-positive elements. This construction takes advantage of the phase transition between zero and near-zero entries, and is closely reminescent of another recent work on so-called apocalypses in the set of rank-bound matrices [10]. This connection suggests explicitly thresholding the elements of the factor matrix $U,L,R$ as a potential safeguard, that might preserve benign nonconvexity by preventing an algorithm from becoming trapped in shallow but combinatorial regions. Developing such safeguarded algorithms or identifying the conditions under which they restore benign nonconvexity could be a fruitful research direction.
Algorithmic implications and approximate stationarity.
In the benign $\delta=0,r^{\star}=1$ regime, we show that all exact constrained second-order stationary points are globally optimal (and that this fails for $\delta>0$ or $r^{\star}>1$). To say that standard algorithms such as gradient flow or gradient descent converge to the global minimum, however, would require the stronger strict saddle property, namely that every $\epsilon$-second-order stationary point lies $O(\epsilon)$-close to a global optimum. Assuming this holds, projected gradient descent (and similar methods) can be equipped with saddle-escape mechanisms to ensure convergence to $\epsilon$-second-order points [14, 11, 15]. On the other hand, in the constrained setting, randomly initialized projected gradient descent can get stuck at a saddle point with positive probability [14]. To fully explain how algorithms actually behave in the benign $\delta=0,r^{\star}=1$ regime, therefore, would require a more refined, problem-specific analysis, which we leave as future work.
Semidefinite programs for constructing counterexamples.
In the analysis of benign nonconvexity in unconstrained low-rank matrix recovery, a key proof technique is the use of semidefinite programming for constructing counterexamples that admit spurious local minima [24, 23, 21]. This is possible because the second-order conditions in the unconstrained case reduce to the linear matrix inequalities $\nabla f(U)=0$ and $\nabla^{2}f(U)\succeq 0$. In the nonnegative setting, the second-order condition requires copositivity $\left\langle\nabla^{2}f(U)[\dot{U}],\dot{U}\right\rangle\geq 0$ over the critical cone $\dot{U}\in\mathscr{C}(U,\nabla f(U))$, which in turn is coupled back to the candidate solution itself, through both $U$ and $\nabla f(U)$. SDP-based constructions do not readily extend, and this is why our counterexample in Theorem 3 was manually constructed by hand. Nevertheless, a more systematic approach to counterexamples could illuminate deeper principles behind why the benign landscape disappears under nonnegativity constraints.
Squared-variable reparameterization
A natural possibility suggested by a reviewer is to eliminate the explicit constraint $U\geq 0$ by reparameterizing $U=V\odot V$ (entrywise), so that the optimization over $V\in\mathbb{R}^{n\times r}$ becomes unconstrained. From the perspective of analyzing benign landscapes, this reparameterization would restore access to the SDP machinery developed for unconstrained problems, as mentioned above. At the same time, it is important to note that the landscape of $\varphi(V)=f(V\odot V)$ can differ qualitatively from the landscape of $f(U)$ over $U\geq 0$. In particular, benignness (or non-benignness) need not transfer between the two parametrizations. Nevertheless, studying basic benign-landscape questions for $\varphi(V)$, and attempting SDP-based counterexample searches in $V$, is a promising direction, as it could yield useful insight into whether alternative parameterizations can mitigate bad landscapes induced by the constraint $U\geq 0$.
Acknowledgments
I am grateful to Xiaohui Chen, Salar Fattahi, Sabrina Zielinski, and three reviewers for helpful feedback and insightful suggestions that significantly strengthened the paper.
Funding
Financial support for this work was provided by NSF CAREER Award ECCS-2047462 and ONR Award N00014-24-1-2671.
Data Availability
Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
Declarations
Conflict of interest
The author has no relevant financial or non-financial interests to disclose.